### v1

# Mission-Aligned Learning-Informed Control of Autonomous Systems: Formulation and Foundations

Abstract
Research, innovation and practical capital investment have been increasing rapidly toward the realization of autonomous physical agents. This includes industrial and service robots, unmanned aerial vehicles, embedded control devices, and a number of other realizations of cybernetic/mechatronic implementations of intelligent autonomous devices. In this paper, we consider a stylized version of robotic care, which would normally involve a two-level Reinforcement Learning procedure that trains a policy for both lower level physical movement decisions as well as higher level conceptual tasks and their sub-components. In order to deliver greater safety and reliability in the system, we present the general formulation of this as a two-level optimization scheme which incorporates control at the lower level, and classical planning at the higher level, integrated with a capacity for learning. This synergistic integration of multiple methodologies—control, classical planning, and RL—presents an opportunity for greater insight for algorithm development, leading to more efficient and reliable performance. Here, the notion of reliability pertains to physical safety and interpretability into an otherwise black box operation of autonomous agents, concerning users and regulators. This work presents the necessary background and general formulation of the optimization framework, detailing each component and its integration with the others.
1 Introduction
Consider a setting in which a robot is a permanent aide to a movement and health-impaired human. This robot is supposed to provide for the individual’s needs, while generally attempting to benefit their health and well-being. It should also ensure that the individual is safe, and that any unexpected behavior should trigger immediately notifying a human specialist with the most critical information. There are various possible relevant objects in the environment, and the robot either moves next to the human to assist movement or retrieves the object for the human. There are a set of regular expected activities, eating, grooming, exercising, scribing dictation, etc. that the robot performs in accordance with: the desires of the human, the human’s well-being as according to the responsible health authorities, the physical safety of the human, and any regulations. The robot can communicate and should suggest any activity on the schedule to the human, who can refuse or stop the activity at any time. Warnings can be issued, e.g., “if we move away now your living room will have this pile and may look messy,” but ultimately the human decides its behavior.
In service of this task, the robot is able to lean to roll forward and back as well as turn in full rotation, and rotate its arm along three axes. Assume, for simplicity, that objects can be picked up without modeling grasping, as if everything relevant was magnetized. Objects in the environment define collisions to be avoided, which requires including the speed and position in the state space. The robot has a propulsion system, which is its forward control, and a braking mechanism. The manufacturing of this propulsion system involves some new features which are not system-identified to full certainty, and there is a small but noticeable nonlinear gyroscopic effect on its motion. This effect will have to be learned in the course of its operation.
This work will present a formalism that defines a bilevel optimization problem incorporating Reinforcement Learning (RL)-assisted Model Predictive Control in the bottom layer, and RL-assisted classical planning at the higher level. In describing the formalism we will return to this example, and in ongoing works analyzing the theoretical properties and testing a simulation of this stylized example.
Human-aligned and safe autonomous robotics is a technology of significant interest to many researchers, companies, and governments, with population aging creating many potential labor shortages for assistance care and health for the elderly, the potential for automated manufacturing, and other reasons. As it is also a representative example of the conceptual components of such an operation, it will be a running example throughout this line of work. In this work we simplify some of these concerns and focus on developing the foundations of the mathematical formulation for the optimization framework. We present a formal schema to define the autonomous decision procedures of an agent using multiple relevant disciplines associated with mathematical engineering.
AI and Autonomous Systems
Designing and managing systems towards optimality or looser satisfaction criteria over time is a generic cross-disciplinary problem that has attracted attention from numerous communities. In the case wherein the model of the states and actions can be well represented by a finite set while the system dynamics and the utility of different circumstances is unknown, Reinforcement Learning (RL) has risen to prominence as the preeminent methodology. RL, like Classification (e.g., image recognition), and Generative Learning (e.g., Large Language Models), has been a central component in the accelerating, groundbreaking developments of AI that have characterized the technological landscape for the past decade. Neural networks leverage the raw computational power of highly parallel GPU hardware permitting training of deep models with millions to billions of trainable parameters. This forms highly accurate statistical models of objects, processes and phenomena when there are vast quantities of data present or available.
In recent years, a variety of approaches for opening the deep and opaque models developed by AI training have appeared aiming to improve their interpretability and reliability. In the case of image recognition and LLMs, these concerns are largely a matter of ensuring an agreeable user experience. However, RL involves making decisions. In the context of games of pure entertainment, this relative lack of criticality of risks persists. However, when RL agents make decisions in contexts with real-world consequences, a notable degree of justified concern appears. An opaque black-box model simply cannot provide formal guarantees of safety or performance reliability. Rather, it is simply a model that efficiently provides a most likely answer, in some formal probabilistic sense. Low probability but catastrophic events could happen.
Contributions
The general premise of this work is as follows: it is currently difficult if not impossible to ensure that RL can systematically avoid or minimize the risk of damaging living beings or critical agents and infrastructure via action or instability. Probabilistic learning techniques, importance sampling in unsafe regions, human in the loop, and other strategies that remain within the RL paradigm garcia2015comprehensive may present clever mechanisms to dramatically lower the probability of said risks, but they are fundamentally insufficient. Comprehensive guarantees of safe and stable operation must arise from a paradigm that is separate from probabilistic inductive learning.
In particular, crisp hard science-based models of system dynamics and foreseeable risks (e.g. of collision) are the standard, and a comprehensively reliable approach to endow RL systems with true guarantees of safe and stable operation. Said guarantees would significantly expand the set of domains and circumstances to which RL can be applied, and at the same time provide a transparent and interpretable means by which regulators can certify the safety of an RL system/agent. This will be done by embedding Model Predictive Control within the lower layer. Moreover, at the highest level of abstraction as to the agent’s operation, the decisions must be interpretable and associated with the mission and tasks associated with the AI. These should make logical sense to a user, and as far as the general agreement and sense of sequences of higher level operations. This will be done by embedding classical planning into the top layer.
In this paper we present a set of technical procedures from the methodological fields of process control and planning that incorporate structural assurances into deep neural networks in a systematic way. Along with addressing the aforementioned concerns of interpretability and reliability, these set of computational techniques significantly improve the quality and precision of the RL model and enable magnitudes of greater statistical power for any available datasets used to train the optimal decision model.
2 Hierarchical Reinforcement Learning
In this Section we provide the background for and the full formulation, as well as a description thereof, of a Bilevel Reinforcement Learning problem as it could be applied to the practical consideration above regarding elder care. To this end, we present a summary of the main principles of RL, its extension to continuous action spaces, and bilevel and generally hierarchical RL problems and solution algorithms.
2.1 Reinforcement Learning - Background
Formally, a standard RL problem (two prominent texts being sutton2018reinforcement ; bertsekas1996neuro ) is defined to consist of a set of states $\mathcal{S}$ and actions $\mathcal{A}$ to find a policy $\pi:\mathcal{X}\to\mathcal{M}(\mathcal{A})$ whereby $\mathcal{M}$ we mean the set of multinomial distributions over the set $\mathcal{A}$. The state evolves in time through an unknown transition probability tensor $\mathbf{P}$ with components $P(S_{t+1}=s^{\prime}|S_{t}=s,A_{t}=a)$. The reward $R(s,s^{\prime},a)$ defines some utility the agent receives upon that particular transition to $s^{\prime}$. In RL, techniques to learn the Value function $V^{\pi}(s)$ and Action-Value function $Q^{\pi}(s,a)$ are iterated together with establishing the policy $\pi$ in either performing the most rewarding action $a$ or randomly explore and learn more about the system.
RL searches for the optimal decision-making policy by interacting with a system modeled as a Markov Decision Process (MDP). Due to its data-driven nature, RL is particularly powerful in complex or stochastic systems, for which explicit models are difficult to construct recht2019tour . RL aims to learn a policy $\pi:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ that minimizes the cumulative reward, or return, defined as:
$\begin{split}\eta(\pi)=\mathbb{E}_{s(0),a(0),\dots}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s(t),a(t))\right],\end{split}$ (1)
where $s(0)\sim\rho(s(0))$, $a(t)\sim\pi(a(t)|s(t))$, and $s(t+1)\sim P(s(t+1)|s(t),a(t))$. At each state, the expected cumulative reward given a policy $\pi$ can be estimated via the value function $V^{\pi}(s(t))$ and the state-action value function $Q^{\pi}(s(t),a(t))$, defined below:
$\begin{split}&V^{\pi}(s(t))=\mathbb{E}_{a(t),s(t+1),\dots}\left[\sum_{i=0}^{\infty}\gamma^{t+i}R(s(t),a(t))\right],\\ &Q^{\pi}(s(t),a(t))=\mathbb{E}_{s(t+1),\dots}\left[\sum_{i=0}^{\infty}\gamma^{t+i}R(s(t),a(t))\right],\end{split}$ (2)
where $a(t)\sim\pi(a(t)|a(t))$, and $s(t+1)\sim f(s(t+1)|s(t),a(t))$. For Eq. 1&2, it is seen that $\eta(\pi)=\mathbb{E}_{s(0)\sim\rho(s(0))}\left[V^{\pi}(s(0))\right]$. With the system modeled as an MDP, state transition depends only on the current state and action, thus, $V^{\pi}(s(t))$ and $Q^{\pi}(s(t),a(t))$ can be recursively estimated from observed interactions $(s(t),a(t),s(t+1),R(s(t),a(t)))$ using the Bellman equations:
$\begin{split}&V^{\pi}(s(t))=\mathbb{E}_{a(t),s(t+1)}\left[R(s(t),a(t))+\gamma V^{\pi}(s(t+1))\right],\\ &Q^{\pi}(s(t),a(t))=\mathbb{E}_{s(t+1)}\left[R(s(t),a(t))+\gamma V^{\pi}(s(t+1))\right],\\ \end{split}$ (3)
where $s(t)\sim\pi(s(t)|a(t))$ and $s(t+1)\sim P(s(t+1)|s(t),a(t))$. The Bellman equations define a recursive update rule for the expected return of state-action pairs, enabling iterative refinement based on observed interactions.
2.2 Reinforcement Learning with Function Approximations
The update rule in Eq. 3 treats states $\mathcal{S}$ and state-action pairs $\mathcal{S}\times\mathcal{A}$ as table entries that need to be estimated. Consequently, methods that directly estimate $V^{\pi}(s(t))$ and $Q^{\pi}(s(t),a(t))$ are referred to as tabular RL. However, in many real-world applications, the state space $\mathcal{S}$ and action space $\mathcal{A}$ are extremely large or continuous, and it becomes infeasible to apply tabular RL due to the curse of dimensionality.
To handle large or continuous state and action spaces, modern RL methods employ function approximation for value functions. Specifically, the policy $\pi$, state-value function $V^{\pi}$, and the action-value function $Q^{\pi}$ are approximated using parameterized functions, $\pi_{\theta}$, $V_{\phi}$, and $Q_{\psi}$, with trainable parameters $\theta$, $\phi$ and $\psi$, respectively. This parameterization facilitates generalization to unobserved states and state-action pairs, allowing learning across $\mathcal{S}$ and $\mathcal{A}$ without explicit enumeration. The function approximation also enables a series of RL algorithms termed the policy gradient methods, which directly optimize the policy $\pi$ to maximize the objective in Eq. 1 sutton1999policy .
Function approximation (in RL), has a number of challenges, such as overfitting and reduced stability during training packer2018assessing ; mnih2013playing . In the tabular setting, where state and action spaces are discrete, each policy update in RL is guaranteed to improve performance watkins1992q ; kakade2002approximately . However, when function approximation is used, such guarantees become difficult to establish. To address stability issues in training RL with deep neural networks as function approximators, various approaches have been proposed. For example, the Replay Buffer mnih2013playing helps stabilize training by sampling batches of historical interactions. Trust-region policy optimization (TRPO) schulman2015trust employs a second-order method to constrain policy updates and prevent performance degradation. Clipped double Q-learning fujimoto2018addressing mitigates overestimation bias in $Q_{\psi}$ learning by leveraging target Q-networks. Still, the performance of RL has been outstanding for many settings - by using deep neural networks as function approximators and the recently proposed stabilizing techniques, RL can now address many problems that were previously intractable levine2016end ; zhu2017target ; degrave2022magnetic .
2.3 Bilevel Reinforcement Learning Formulation
Hierarchical, or Multilevel (in generalizing the classic Bilevel case) optimization concerns the solution of problems whereby the criteria exhibit a dependent layer structure. That is, the reward function for the top layer depends on the value of another variable which is in turn the solution of a different optimization, one which in turn depends on the operations of the earlier decision variables. See, e.g. migdalas2013multilevel for a collection of representative works. When all of the levels define convex optimization problems, reliable single loop algorithms can be assured through careful step level-weighing strategies shafiei2024trilevel .
In the case wherein the cost/reward functions are nonconvex, convergence guarantees are less prevalent and firm. Standard Stochastic Approximation multilevel extensions are described in dereich2019general . Hierarchical Stochastic Approximation has also been considered for (nonsmooth) risk measures crepey2024adaptive . Typically, these algorithms require a careful tuning of stepsize annealing, with a hierarchy of decay rates for stepsizes associated with each level. However, see a recent single-timescale analysis in ghadimi2020single and shen2022single .
However, in consideration of Reinforcement Learning, there is an additional bilevel structure at each layer. In particular, the policy is what is of interest, but the policy depends on the Value function associated with each state, which itself must be learned. Since this dependency is directional, two timescales is the standard approach, e.g. zeng2024two . For an application in multi-stage stochastic programming, which can be considered an approximation to DP, with Stochastic Approximation, see lan2021dynamic . The two timescale approach lends itself to theoretical analysis of the learning procedure using Stochastic Differential Equations, due to the work in e.g. yin2012continuous ; zhang2017optimal ; yin2009asymptotic ; nguyen2024second . Extensions to the hierarchical case, that is, considering both the hierarchy of DP problems together with the necessity of two time scale methods for RL, will be necessary for implementing this work.
For Reinforcement Learning specifically kreidieh2019inter considers a boss-worker hierarchical RL with derivation of a cooperative policy gradient for the boss. The work zhu2022hierarchical defines navigation to avoid obstacles at a lower level, and goal-driven movement at the upper level, for a robotics RL engine. Image captioning with a language layer and a visual layer of RL is described in xu2019multi
To define the two tiered problem of managing a robotic aid to a human, we use a discrete action space for the set of chosen tasks, such as to clean, feed, etc. for the upper problem and a continuous action space for the optimal control defining navigation and operation at the lower level. We will use deterministic policies, meaning that there is one unique optimal action chosen upon the presence of any given state. Of course, algorithmically, we may still randomize the action chosen for exploration purposes.
The Upper Level RL Problem aims to learn a deterministic policy $\pi:\mathcal{X}\to\mathcal{A}$ that minimizes the cumulative stage cost defined as:
$\eta(\pi)=\mathbb{E}_{x(0,x(1),\cdots)}\left[\sum_{t=0}^{\infty}\gamma^{t}F\left(x(t),a(t),\left\{z(\tau),v(\tau\}\right\}_{t-1<\tau\leq t}(x(t-1),a(t-1))\right)\right],$
where
$x(0)\sim\rho(x(0)),a(t)=\pi(s(t))$
and
$x(t+1)\sim f(x(t+1)|x(t),a(t),\{z(\tau),v(\tau)\}_{t<\tau\leq t+1})$
and $z(\tau)$ is a time varying continuous state space encoding the location, velocity and other canonical coordinates. Note that this $f$ encodes a probability distribution over a set of states $x(t+1)\in\mathcal{X}$. The inner problem control $v(\tau)$ is the time varying control action driving a deterministic process $z(\tau+h)=g(z(\tau),v(\tau))$ where $h\ll 1$ is the fast time scale. Note that this process depends on the state and action of the upper level chosen at the previous time step.
The continuous control action $v(\tau)$ solves a lower level problem:
$\begin{split}\zeta(\pi;\eta,t)=\mathbb{E}_{x(t+1)}\left[\sum_{\tau=t,t+h,\cdots}^{t+1}\beta^{\tau}G\left(z(\tau),v(\tau),x(t+1),a(t+1),x(t),a(t))\right)\right],\end{split}$ (4)
where $G$ is a functional that includes either $F$ and other local physical criteria, or tracking a solution that optimizes the top level functional.
In this case the upper level value function $V^{\pi}(x(t))$ becomes:
$\begin{split}V^{\pi}(x(t))=\mathbb{E}_{x(0),x(1),\cdots}\left[\sum_{i=0}^{\infty}\gamma^{t+i}F(x(t),a(t),\left\{z(\tau),v(\tau\}\right\}_{t-1<\tau\leq t}(x(t-1),a(t-1)))\right],\\ a(t)=\pi(x(t))\end{split}$ (5)
However, this would assume $\left\{z(\tau),v(\tau\}\right\}_{t-1<\tau\leq t}(x(t-1),a(t-1))$ already optimizes the Value function, that is satisfies the Bellman equations
$\begin{split}V^{\zeta}(z(\tau))=\mathbb{E}_{x(t+1)}\left[\beta^{\tau}G\left(z(\tau),v(\tau),x(t),a(t),x(t+1),a(t+1))\right)+\beta V^{\zeta}(z(\tau+h))\right],\end{split}$ (6)
for all $t$. Observe the inherent coupling, the upper and lower level solutions depend on each other, introducing a couple Dynamic Programming Fixed Point Problem. However, one can consider that in real-time operation, one can predict the state for future control problems, and so always have precomputed solutions one is consistently updating.
3 Safe RL with Model Predictive Control
A significant challenge in applying RL to real-world problems is safety. Although approaches based on Constrained Markov Decision Processes (CMDPs) altman2021constrained ; achiam2017constrained allow minimization of the expected cumulative stage cost while following a set of safety constraints, they often require the exploration of unsafe trajectories during training. To avoid safety violations during training in particular (and to enhance learning efficiency in general), an emerging solution is the integration of domain knowledge or system constraints into the RL learning process garcia2015comprehensive .
The relevant domain for ensuring safe operation according to the laws of physics is Model Predictive Control (MPC). Indeed, the primary historical distinction between the two fields, model predictive control and reinforcement learning, developed under their separate domain native considerations, is the use of explicit background knowledge informed from known scientific principles, on the one hand, and (often very) large sample experiments querying a black box comprehensive roll-out simulation of the entire procedure. In practice, one is often in between the ideal scenario for MPC and R, that is, a partially complete physical understanding of some of the mechanics is available. As such, part of the system dynamics are known, and part have to be ascertained with data.
3.1 Model Predictive Control - Background
Formally, MPC consists of repeated solution of an optimal control problem (OCP) defined over some finite horizon as,
$\begin{array}[]{rl}x^{*},u^{*}\in&\arg\min\limits_{x(t)\in\mathbb{R}^{x},u(t)\in\mathcal{C}^{u}}\frac{1}{T}\sum\limits_{t=1}^{T}F(x(t),u(t))\\ \text{subject to }&x(t)=f(u(t),x(t-1))\\ &c^{I}(u(t),x(t))\leq 0,\,\forall t\\ &c^{E}(u(t),x(t))=0,\,\forall t\\ \end{array}$ (7)
where $f$ is a discretized differential equation (there are variants with implicit integrators as well, we simplify for presentation), and $c$ are constraints. There are control $u$ and state $x$ variables and $\mathcal{C}^{u}$ is some simple (e.g. bounds) constraint set
With MPC, one solves this problem, applies the first control, lets the system evolve, measures the state again, and solves a new OCP with the horizon shifted one step forward. Note two important distinctions from RL: 1) the lack of uncertain quantities or random variables in the model, and 2) the presence of functions defining constraints. These constraints can enforce safety, e.g., by indicating the geometry of the environment as far the presence of objects an agent can collide with.
MPC provides a policy by solving an OCP at each discrete time instant based on the current system state $s$, on a finite receding horizon. The MPC problem can be cast as,
$\min_{x,u}$ $\quad V_{T}\left(x_{T}\right)+\sum_{t=0}^{T-1}\,F\left(x_{t},u_{t}\right)\,,$ (8a)
$\mathrm{s.t.}$ $\quad x_{t+1}=f\left(x_{t},u_{t}\right),\quad x_{0}=s\,,$ (8b)
$\quad h\left(x_{t},u_{t}\right)\leq 0,\quad u_{t}\in\mathcal{A}\,,\quad$ (8c)
where $T$ is the prediction horizon, $F$ is the stage cost, $V_{T}$ is the terminal cost, $f$ is the dynamics and $h$ is the inequality constraint. Problem 8 produces a complete profile of control inputs $u^{\star}=\{u_{0}^{\star},\ldots,u_{N-1}^{\star}\}$ and corresponding state predictions $x^{\star}=\{x_{0}^{\star},\ldots,x_{N}^{\star}\}$. However, only the first element $u_{0}^{\star}$ of the input sequence $u^{\star}$ is applied to the system MPCbook . At the next physical sampling time, a new state $s$ is received, and problem 8 is solved again, producing a new $u^{\star}$ and a new $u_{0}^{\star}$. MPC (8) hence yields a policy
$\pi_{\mathrm{MPC}}\left(s\right)=u_{0}^{\star}\,,$ (9)
with $u_{0}^{\star}$ solution of 8 for $s$ given. For $\gamma\approx 1$, policy (9) can provide a good approximation of the optimal policy $\pi^{\star}$ for an adequate choice of prediction horizon $T$, terminal cost $V_{N}$ and if the MPC model $f$ approximates the true dynamics sufficiently well.
3.2 Model Predictive Control and RL as Dynamic Programming
Ultimately, one can consider both model predictive control and reinforcement learning to be problems that simplify the general Dynamic Programming, or Stochastic Control, problem. This resemblance and the ultimate possibility of unifying analysis of algorithmic structure in these historically distinct research communities was pointed out in powell2014clearing ; powell2019unified .
$\inf\limits_{u(t)\in\mathcal{C}^{u}\subseteq\mathbb{R}^{n_{u}},\forall t}$ $\mathbb{E}_{\xi}\bigg{[}\sum_{t=1}^{T}F(x(t,\xi(1:t)),u(t,\xi(1:t)),\xi(t))\bigg{]}$ (10)
subject to $x(1,\xi(1))=f(u(1,\xi(1)),x_{0},\xi(1)),$
$x(t,\xi(1:t))=f(u(t,\xi(1:t)),x(t-1,\xi(1:t-1)),\xi(t)),\quad t=2,\dots,T,$
$c^{I}(u(t,\xi(1:t)),x(t,\xi(1:t)))\leq 0,\quad t=1,\dots,T,$
$c^{E}(u(t,\xi(1:t)),x(t,\xi(1:t)))=0,\quad t=1,\dots,T.$
Stochastic MPC can be considered a look-forward policy approximated with SAA. RL can be considered developing an empirical model wherein the constraints and the dynamics are completely encoded as embedded in the deep representation of the Value/Action-Value function. Thus, it can be seen that both schemes are different means of approximating a dynamic programming problem. In MPC, a continuous-time control-state system is discretized, and an iterated sequence of lookahead problems is solved. In RL, trajectory sampling as information gathering is a central component of the algorithmic process, however only problems with finite state and action spaces become tractable. This choice of approximation/modeling limitation comes from the domain, and for our purposes, in order to ensure physical constraints on the system, we will use discrete time but continuous state-action space representations . However, rather than being some latent variable or structure in the value function, the influence of the actions and states are now subject to nonlinear equation systems. Thus domain physics knowledge constrains the continuous space sufficiently so as to become computationally tractable, at least with approximate techniques.
3.3 Model Predictive Control Incorporation within RL
In recent years the integration of Reinforcement Learning (RL) with Model Predictive Control (MPC), known as RL-based MPC (RLMPC), has gained significant attention for addressing the limitations of conventional MPC methods in dynamic and uncertain environments. This approach leverages MPC as a function approximator within RL kordabad2023reinforcement and has been proven to yield optimal solutions to Markov Decision Processes (MDPs) even when the predictive model is inaccurate gros2019data . The RLMPC framework proposed in gros2019data combines the strengths of both paradigms: MPC ensures explainable policies with guarantees, while RL improves closed-loop performance of MPC by adapting it directly from data to compensate for model inaccuracies. Furthermore, RLMPC can encode explicit safety constraints to provide a safe-RL framework gros2022learning ; zanon2020safe . These advancements position RLMPC as a powerful framework for handling complex, constrained, and adaptive decision-making problems across various applications.
RLMPC approaches have been increasingly applied to real-world systems, especially in energy systems and robotics. In energy systems, RLMPC has been applied to optimize peak power management and energy flexibility in smart grids, including multi-agent residential systems with local renewable energy production and storage cai2021optimal ; kordabad2021multi ; cai2023learning . In robotics, RLMPC has been used for autonomous surface vessels (ASVs) to enhance their performance in tasks such as collision avoidance, docking, and trajectory tracking kordabad2021reinforcement ; cai2021mpc ; martinsen2022reinforcement .
The performance of an MPC scheme relies heavily on the accuracy of its predictive model, yet many real-world systems are inherently difficult to model precisely anand2024optimality ; anand2024data . This model dependency is a key limitation of MPC, as selecting the model that maximized the closed-loop performance is non-trivial, and a more accurate model does not always guarantee better performance. However, it has been shown that MPC can still achieve optimality under an inaccurate model by appropriately adjusting its parameters gros2019data . The integration of Reinforcement Learning (RL) with MPC aims to overcome this limitation by using RL to adaptively improve the closed-loop performance. Recent studies have established that when focusing on closed-loop performance, an MPC scheme can be adjusted holistically, allowing for the design of optimal policies without relying on a completely accurate model of the system. As another strategy, we can also see the use of explicit information learning methods from Partially Observed MDPs applied to learning an uncertain model in matias2022simultaneous .
Now we will introduce the necessary background on RLMPC. Consider an MPC-based policy,
$\pi_{{\theta}}(s)=u^{\star}_{0}\,,$ (11)
where $u^{\star}_{0}$ is part of the solution of:
$x^{\star},u^{\star}=\mathrm{arg}\min_{x,u}$ $\quad V_{{\theta}}\left(x_{N}\right)+\sum_{k=0}^{N-1}\,F_{{\theta}}\left(x_{k},u_{k}\right),$ (12a)
$\mathrm{s.t.}$ $\quad x_{k+1}=f_{{\theta}}\left(x_{k},u_{k}\right),\quad x_{0}=s\,,$ (12b)
$\quad h_{{\theta}}\left(x_{k},u_{k}\right)\leq 0,\quad u_{k}\in\mathcal{A}\,.$ (12c)
This MPC formulation is identical to (8), but the cost, constraints, and dynamics underlying the MPC scheme are now all parameterized in ${\theta}$, to the exception of the input constraint $u_{k}\in\mathcal{A}$. Besides the MPC-based policy (11), MPC can be construed as a (possibly local) model of the action-value function $Q^{\star}$. We can define a parameterized action-value function $Q_{{\theta}}$ based on the MPC scheme (12), as a model of the optimal action-value function $Q^{\star}$ as follows:
[…]
where a constraint $u_{0}=a$ included in (13b) is the only difference to (12). MPC (13) is a valid model of $Q^{\star}$ in the sense that it satisfies the following Bellman relationships on the true system
$V^{\star}\left(s\right)$ $=\min_{a}\,Q^{\star}\left(s,a\right)\,,$ (14a)
$Q^{\star}\left(s,a\right)$ $=F\left(s,a\right)+\gamma\mathbb{E}\left[V^{\star}\left(s_{+}\right)\,|\,s,a\,\right]\,.$ (14b)
and the optimal policy
$\pi^{\star}\left(s\right)=\mathrm{arg}\min_{a}\,Q^{\star}\left(s,a\right)\,.$ (15)
i.e.:
$V_{{\theta}}(s)$ $=\min_{a}\,Q_{{\theta}}(s,a)\,,$ (16a)
$\pi_{{\theta}}\left(s\right)$ $=\mathrm{arg}\min_{a}Q_{{\theta}}(s,a),$ (16b)
where $V_{{\theta}}(s)$ is the optimal cost resulting from solving MPC (12). If the MPC parameters $\theta$ are such that $Q_{{\theta}}=Q^{\star}$, then the MPC scheme (12) delivers the optimal policy $\pi^{\star}$ through (11), i.e. $\pi_{{\theta}}=\pi^{\star}$. An important question, then, is how effectively an MPC scheme can approximate $Q^{\star}$ at least in a neighborhood of $a=\pi^{\star}\left(s\right)$. The central result in RL-based MPC provided in Theorem 1 in gros2019data answers this question and establishes it under some mild conditions, (13) provides an exact model of $Q^{\star}$ even if the predictive model (12b) is inaccurate. The Theorem 1 in gros2019data states the following:
Considering the parameterized stage cost, terminal cost, and constraints in (12) are universal function approximations (i.e., can approximate a given function accurately) with adjustable parameters $\theta$. Then there exist parameters $\theta^{\star}$ s.t. the following identities hold $\forall\gamma$:
  1. 1.
    $V_{\theta^{\star}}(s)=V^{\star}(s),\,\forall s\in\Omega$
  2. 2.
    ${\pi}_{\theta^{\star}}(s)=\pi^{\star}(s),\,\forall s\in\Omega$
  3. 3.
    $Q_{\theta^{\star}}(s,a)=Q^{\star}(s,a),\,\forall s\in\Omega$, for the inputs $a\in\mathcal{A}$ such that $\lvert V^{\star}(f_{\theta^{\star}}(s,a))\rvert<\infty$
if the set
$\Omega$ =$=\left\{s\in\mathcal{S}\,\,\Big{|}\,\,\left|V^{\star}(x^{\star}_{t})\right|<\infty,\ \forall\,t\leq T\right\}\,,$ (17)
is non-empty, where $x^{\star}_{t}$ is the optimal state solution of (12).
This in turn entails that MPC (12) can provide the optimal solution to the true MDP even if the MPC model is inaccurate. This can be achieved by selecting the appropriate stage cost, terminal cost, and constraints in the MPC scheme. The applicability of this results is very broad and extends, e.g., to robust MPC, stochastic MPC, mixed-integer MPC, and Economic MPC (EMPC), and is valid both for the discounted and undiscounted settings zanon2020safe ; kordabad2022safe ; gros2020reinforcement . The role of RL in that context is to provide these learning tools . RL-based MPC integrate RL tools that aim at minimizing $J(\pi_{\theta})$, or at achieving $Q_{{\theta}}\approx Q^{\star}$ from the data. There are alternatives to RL for adjusting the MPC parameters, such as e.g. Bayesian optimization. However, RL is often regarded as the most effective technique, especially if the number of parameters to adjust is not very small.
Both Q-learning and policy gradient approaches can be used for RL-based MPC kordabad2023reinforcement ; zanon2019practical . In case of Q learning aim to satisfy (3), where $Q_{\theta}$ is delivered by (13) and $\min_{a^{\prime}}Q_{\theta}(s,a^{\prime})=V_{{\theta}}(s)$ is delivered by (16b), i.e. by the optimal cost of MPC (12) solved at state $s$. Alternatively policy gradient approach aim to adjust the parameters $\theta$ of the MPC as a policy (11) to minimize cumulative cost $J(\pi_{\theta})$ directly to satisfy (2). This is the standard approach for systems with continuous action spaces for which we can distinguish between two methods. Policy gradient methods estimate the policy gradient $\nabla_{\theta}J(\pi_{\theta})$ from data to update the parameters $\theta$ in a gradient descent fashion. Alternatively, direct policy search methods build a surrogate model of $J(\pi_{\theta})$ from data and use it to propose new policy parameters $\theta$. However, direct policy search tends to scale poorly with the size of the policy parameters.
One common approach in policy gradient methods is the deterministic policy gradient by silver2014deterministic with the following expression for the gradient of the cumulative cost:
$\nabla_{\theta}J(\pi_{\theta})=\mathbb{E}\left[\nabla_{\theta}\pi_{\theta}\left(s\right)\nabla_{a}Q_{\pi_{\theta}}\left(s,\pi_{\theta}(s)\right)\right]\,.$ (18)
Here $\pi_{\theta}$ is delivered by MPC (12). The critic $Q_{\pi_{\theta}}$ is typically built separately using a generic function approximator and policy evaluation techniques. It has been shown that the DPG method can be simplified in the case of RLMPC by approximating the required value function through a first-order expansion of the value function estimate obtained from the MPC policy itself anand2023painless .
3.3.1 Sensitivities
Many RL methods, including Q-learning and policy gradient methods, require the sensitivities of the function approximators $\pi_{\theta}$, $Q_{\theta}$, and $V_{\theta}$. When approximated by an MPC scheme, these are typically continuous but only piecewise smooth. However, when the MPC achieves Linear Independence Constraint Qualification (LICQ) and Second Order Sufficient Condition (SOSC), non-smooth points correspond to weakly active constraints in the MPC. These points form a set of zero measures for a well-formulated MPC scheme. Because RL methods always use the sensitivities inside expected value operators, their contribution to the learning disappears as long as the state transition density of the true MDP is bounded. The non-smoothness of MPC schemes may superficially appear as an issue for RL, but fortunately, in all practical cases, it is not. This issue can be ignored when the MPC response is continuous.
The arguments above do not necessarily hold anymore if the MPC does not satisfy SOSC and, as a result, produces discontinuous policies, e.g., a bang-bang response. For example, designing the MPC as a Linear Program might lead to such a situation, as well as processes with switching operations or more decisions than state equations. We can then still use Q-learning since $Q_{\theta}$ and $V_{\theta}$ typically remain continuous and piecewise smooth. However, the discontinuities in the policy lead to problems in policy gradient methods. An early approach to solve this issue is formulated in kordabad2021mpc , but more work is necessary. In addition, extending optimization methods that perform MPC for problems with degenerate constraints, those that do not satisfy the LICQ, like kungurtsev2017predictor , to incorporate RL, remains an important open problem.
3.3.2 Real-time MPC
An MPC problem at a give initial state $s$ is a Nonlinear Program of the form
$\min_{w}$ $\quad\Phi(w)$ (19a)
$\mathrm{s.t.}$ $\quad g(w,s)=0$ (19b)
$\quad h(w)\leq 0$ (19c)
where $w$ gathers the discrete-time state and input trajectories. The equality constraints (19b) typically ensure that the state trajectories abide by the physical model (8b) used in formulating the MPC problem.
Defining the Lagrange function associated to (19):
$\mathcal{L}(w,\lambda,\mu)=\Phi(w)+\lambda^{\top}g(w,s)+\mu^{\top}h(w)$ (20)
for the dual variables $\lambda,\mu$, a solution to NLP (19) satisfies the KKT conditions:
$\nabla_{w}\mathcal{L}(w,\lambda,\mu)=0$ (21a)
$g(w,s)=0$ (21b)
$h(w)\leq 0,\quad\mu\geq 0$ (21c)
$\mu^{\top}h(w)=0$ (21d)
for dual variables $\lambda,\mu$. If NLP (19) holds a sufficiently differentiable cost $\Phi$ and constraints $f,h$, it can be solved via Successive Quadratic Programming (SQP), whereby problem (19) is successively approximated by the Quadratic Program (QP)
$\min_{\Delta w}$ $\quad\frac{1}{2}\Delta w^{\top}H\Delta w+\nabla_{w}\Phi(w)^{\top}\Delta w$ (22a)
$\mathrm{s.t.}$ $\quad g(w,s)+\nabla_{w}g(w,s)^{\top}\Delta w=0$ (22b)
$\quad h(w)+\nabla_{w}h(w)^{\top}\Delta w\leq 0$ (22c)
where matrix $H$ is the Hessian with respect to $w$ of the Lagrange function $\mathcal{L}$ at $w,\lambda,\mu$. The solution of QP (22) provides the primal-dual update:
$w\leftarrow w+\alpha\Delta w,\quad\lambda\leftarrow\lambda+\alpha\Delta\lambda,\quad\lambda\leftarrow\mu+\alpha\Delta\mu$ (23)
where $\Delta\lambda$ and $\Delta\mu$ are given by the change in the dual variables associated to (22) from the previous QP solution, and $\alpha\in(0,1]$ is a step size selected via a line-search strategy. SQP iterates solving QP (22) and updating the primal-dual variables (23) until convergence to a solution $w^{\star}(s),\lambda^{\star}(s),\mu^{\star}(s)$ to (21) for a given $s$. The SQP procedure start with an initial guess $w_{\mathrm{init}}$ at which the first QP (22) is formed. If $w^{\mathrm{init}}$ is sufficiently close to $w^{\star}$, then full steps $\alpha=1$ can be used in (23), and the SQP iteration converges quadratically to $w^{\star}(s)$.
A key observation for real-time MPC is that the initial condition $s_{k}$ at a given discrete time step $k$ are typically close to their prediction formed at time $k-1$. More specifically, considering the predicted optimal state $x_{1}^{\star}(s_{k-1})\subset w^{\star}(s_{k-1})$ at time $k-1$ is typically a good approximation of $s_{k}$. This observation leads to the shifting strategy to provide an initial guess for NLP (19) at $s_{k}$ from the solution obtained at time $k-1$ for $s_{k-1}$. The shifting strategy consists in using the optimal state-input trajectories $x^{\star}_{0,\ldots,N}(s_{k-1})$ $u^{\star}_{0,\ldots,N-1}(s_{k-1})$ contained in $w^{\star}(s_{k})$ to form the state-input initial guesses:
$x^{\mathrm{init}}_{i}(s_{k})$ $=x^{\star}_{i+1}(s_{k-1}),\quad u^{\mathrm{init}}_{i}(s_{k})$ $=u^{\star}_{i+1}(s_{k-1}),\quad i=0,\ldots,N-1$ (24)
The new state and input $x^{\mathrm{init}}_{N}(s_{k})$, $u^{\mathrm{init}}_{N-1}(s_{k})$ cannot be built from (24), and are often simply guessed using
$u^{\mathrm{init}}_{N-1}(s_{k})$ $=u^{\star}_{N-2}(s_{k-1}),$ (25a)
$x^{\mathrm{init}}_{N}(s_{k})$ $=f(x^{\star}_{N-1}(s_{k-1}),u^{\star}_{N-2}(s_{k-1}))$ (25b)
More advance strategies are possible but do not always make a significant difference in the performance of the SQP iteration. Initial guesses for the dual variables $\lambda,\mu$ can also be built using a similar process.
A key to ensuring that $x_{1}^{\star}(s_{k-1})\approx s_{k}$ is to make sure that the sampling time of the MPC scheme is fast enough to ensure that the prediction error cannot accumulate during the physical time interval between the discrete times $k-1$ and $k$. When this strategy is sufficiently effective, the SQP procedure using the initial guess (24) often fully converges in a very few iterations. An extreme version of SQP for real-time MPC is then to perform a single SQP iteration at each time sample $k$ of MPC, and rely on the strong quadratic contraction resulting from the update (23) from the initial guess (24)-(25) to keep the (not fully converged) primal-dual solutions close to their optimal values $w^{\star}(s),\lambda^{\star}(s),\mu^{\star}(s)$. This principle is at the core of the Real-Time Iteration (RTI). In addition to circumventing multiple SQP iterations, RTI seeks to minimize the time spent between obtaining a state measurement $s_{k}$ and delivering the corresponding approximation to the optimal solution $w^{\star}(s_{k})$. Indeed, delivering $w^{\star}(s_{k})$ requires performing two computationally expensive operations: forming QP (22), i.e. computing $H,\nabla_{w}\Phi(w),g(w,s_{k}),\nabla_{w}g(w,s),h(w),\nabla_{w}h(w)$, and solving it. A key observation behind the RTI is that it is possible to form QP (22) before receiving $s_{k}$. Hence the RTI performs the update of the MPC solution in two phases. The preparation phase consists in performing the shifting (24)-(25) and preparing QP (22), while the feedback phase consists in solving QP (22) once $s_{k}$ is received.
3.3.3 Exploration in RLMPC
Exploration is a core requirement for all RL methods, i.e., regularly applying actions $\mathbf{a}\neq\mathbf{\pi}_{\mathbf{\theta}}(\mathbf{s})$ to the real system to gather the information necessary to improve the policy. Given the constraints in (12c), a natural question is how to generate exploration that does not jeopardize the MPC feasibility. A straightforward way to address this is by manipulating the MPC cost such that the resulting action differs from the original policy provided by (12). One possible modification is to add a gradient over the initial action $\mathbf{u}_{0}$, which yields the MPC scheme
$\phi_{\mathbf{\theta}}\left(\mathbf{s},\mathbf{d}\right)=\min_{\mathbf{x},\mathbf{u}}$ $\,\mathbf{d}^{\top}\mathbf{u}_{0}+T_{\mathbf{\theta}}\left(\mathbf{x}_{N}\right)+\sum_{k=0}^{N-1}\,L_{\mathbf{\theta}}\left(\mathbf{x}_{k},\mathbf{u}_{k}\right)$ (26a)
[…]
where $\mathbf{d}\in\mathbb{R}^{m}$ is a vector of the size of the action, possibly selected randomly. Because only the MPC cost is modified, the solution of (28) is feasible for (12). One can then use MPC (28) to produce feasible exploration for policy gradient methods and to generate action-value functions for Q-learning. In that context, (28) produces a feasible action with exploration $\mathbf{a}=\mathbf{u}_{0}^{\star}$ where $\mathbf{u}_{0}^{\star}$ is the solution of (28) and depends on $\mathbf{d}$. The optimal cost $\phi_{\mathbf{\theta}}$ of (28) delivers the action-value function
$Q_{\mathbf{\theta}}\left(\mathbf{s},\mathbf{a}\right)=\phi_{\mathbf{\theta}}\left(\mathbf{s},\mathbf{d}\right)-\mathbf{d}^{\top}\mathbf{u}_{0}^{\star}\,.$ (27)
see gros2019data for more details. However, the feasibility of (28) does not guarantee that applying the resulting policy to the real system will not violate the constraints (12c). For example, consider the effect of stochastic dynamics or model errors. This issue of safe exploration is discussed in the next section.
An additional technique is presented in matias2022simultaneous . Here, one solves the modified problem:
$\phi_{\mathbf{\theta}}\left(\mathbf{s},\mathbf{d}\right)=\min_{\mathbf{x},\mathbf{u}}$ $\,T_{\mathbf{\theta}}\left(\mathbf{x}_{N}\right)+\sum_{k=0}^{N-1}\,\left\{L_{\mathbf{\theta}}\left(\mathbf{x}_{k},\mathbf{u}_{k}\right)+JS(\mathbf{x}_{k};Q)\right\}$ (28a)
[…]
where $JS(\mathbf{x}_{k};Q)$ is the Jensen-Shannon divergence, an information metric, associated with the addition information provided to $Q$ upon potential visit to state $\mathbf{x}_{k}$. This includes the information content as by the reduction in uncertainty of the RL functions explicitly.
3.3.4 Data Drive Predictive Control
In scenarios where explicit state-space models are unavailable, recent advancements have extended the framework of RLMPC to operate without predefined notions of state or explicit dynamics models sawant2023model . This broadens the applicability of RLMPC to systems with partially or completely unknown dynamics. Similarly, other studies have introduced end-to-end learning approaches for Nonlinear MPC (NMPC) policies, enabling the derivation of control strategies directly from data without relying on first-principles models adhau2024reinforcement . These methods are particularly relevant in settings where deriving accurate system models is impractical or infeasible.
Computational complexity is another major concern when combining RL with MPC, especially when gradient-based methods are used. These methods often require frequent evaluations of computationally expensive MPC schemes, limiting their scalability to large datasets. To address this, approaches have been proposed to learn parameterized MPC schemes directly from offline data, bypassing the need to solve MPC problems repeatedly over the collected dataset sawant2023learning . Broadly speaking, however, this is an important open area of research.
3.3.5 Safe Reinforcement Learning with RLMPC
The research community is increasingly considering the application of RL to safety-critical systems. A straightforward way to define safety is through a set of critical constraints which should not be violated, typically expressed as functions of the system state. A classic approach to safe RL is to learn the policy in silico and employ a pessimistic model of the real system, meaning that it overestimates the probability of violating critical constraints. During learning, one can then use high penalties for violations of the critical constraints, e.g. through the use of barrier functions anand2021safe . RL will then naturally adjust the policy to avoid these penalties.
In MPC, using a pessimistic model of the system is common within the Robust MPC (RMPC) methods. Starting from a pessimistic model, RMPC builds a safe policy by ensuring that even the worst-case predictions satisfy the critical constraints at all future times. To that end, the deterministic model (12b) is replaced by a model that describes the evolution of sets enclosing all possible future trajectories, leading to an expression of the form
$\mathbb{X}_{k+1}=\mathbf{f}_{\mathbf{\theta}}(\mathbb{X}_{k},\mathbf{\pi}_{\mathrm{c}}(\mathbb{X}_{k},\mathbf{x}_{k},\mathbf{u}_{k}))\oplus\mathbb{W}_{\mathbf{\theta}}\,.$ (29)
One then designs the policy $\mathbf{\pi}_{\mathrm{c}}$ to manage the growth of the sets $\mathbb{X}_{k}$, usually by operating on their deviation from a reference trajectory $\mathbf{x}_{k}$. The set $\mathbb{W}_{\mathbf{\theta}}$ represents process noise to capture the prediction uncertainties. It is possible to learn (29) from data through set-membership identification. We can then build the following MPC scheme to enforce satisfaction of constraints (12c) explicitly:
$\min_{\mathbf{x},\mathbf{u},\mathbb{X}}$ $\quad T_{\mathbf{\theta}}\left(\mathbf{x}_{N}\right)+\sum_{k=0}^{N-1}\,L_{\mathbf{\theta}}\left(\mathbf{x}_{k},\mathbf{u}_{k}\right)$ (30a)
[…]
$\quad\mathbf{h}_{\mathbf{\theta}}\left(\mathbb{X}_{k},\mathbf{\pi}_{\mathrm{c}}(\mathbb{X}_{k},\mathbf{x}_{k},\mathbf{u}_{k})\right)\leq 0,$ (30c)
$\quad\mathbf{\pi}_{\mathrm{c}}(\mathbb{X}_{k},\mathbf{x}_{k},\mathbf{u}_{k})\in\mathcal{A}$ (30d)
$\quad\mathbb{X}_{N}\in\mathbb{T}_{\mathbf{\theta}},\quad\mathbf{x}_{0}=\mathbf{s},\quad\mathbb{X}_{0}=\mathbf{s}\,.$ (30e)
MPC (30) generates a safe policy $\mathbf{\pi}_{\mathbf{\theta}}=\mathbf{u}^{\star}_{0}$ by construction for $\mathbb{T}_{\mathbf{\theta}}$ adequately chosen and if (29) accounts for the worst case situations observed in the data, see MPCbook . Application of RL to adjust RMPC schemes has been proposed in zanon2020safe . In the context of RL for RMPC-based policies, it is useful to stress that there is a hard separation between learning for safety and learning for closed-loop performance. Indeed, safety is learned via set-membership identification on (29) and then enforced in the RMPC scheme by construction. Closed-loop performance is optimized using RL in parallel (zanon2020safe, ).
For the sake of clarity, we ought to underline that the adjustment of the constraints (30c) in the RMPC scheme and of set $\mathbb{W}_{\mathbf{\theta}}$ ought to be done with care in order to preserve safety. In particular, the adjustment of set $\mathbb{W}_{\mathbf{\theta}}$ must ensure that model (29) accounts for all past data points in the set-membership sense. Arguably, safety-critical constraints in (30c) ought not to be modified.
Safe Exploration
Safe exploration is difficult to produce without a model of the system in the form (29), which can predict the worst-case evolution of the system, and assess the impact of the exploration on the system safety. However, even with a model (29) of the system, it can be expensive to verify the safety of an input differing from the safe policy, and even more expensive to build the set of safe inputs.
It is straightforward to apply the feasible exploration approach discussed earlier to the RMPC formulation (30). Given that we only modify the RMPC cost, the resulting solution is feasible for (30), therefore enabling safe exploration if (30) yields a safe policy. This principle has been further detailed in zanon2020safe .
Safe Policies and Safe Learning
There are two ways to implement safe learning online. One is in a batch fashion, meaning that parameter updates require collecting a minimum amount of transition data. The alternative is to perform parameter updates at every time step. In safety-critical applications, safety must always be preserved in either case.
However, while taking actions from a unique, safe, and stable policy ensures the stability and safety of the system, taking actions from a sequence of safe and stable policies may not. That is because a sequence of policies does not necessarily inherit the properties of the individual policies. We discuss these conditions in detail in gros2022learning .
3.4 Soft Constraints
In many real-world safe control applications, the “estimated” environment models are available (or can be built) to derive sub-optimal control policy priors liepe2014framework ; hippisley2017development . Control policy priors of reasonable accuracy are essential to many applications for critical systems where control failure is consequential, especially systems with a single-life setting where failure even during training is unacceptable. Consider the example of devising a control policy that prescribes the optimal drug dosages for regulating a patient’s health status. This is a critical system with a single-life setting where no harm to the patient is tolerated during policy exploration. From available records of other patients, an estimated patient model can be built to predict the response to different drug dosages and ensure adherence to the safety bounds (set based on clinical knowledge). However, a new patient’s response can deviate from the estimated model, which poses a significant challenge in control adaptability and patient treatment performance.
In applications analogous to the above example, i.e., control in critical systems, MPC derived from the estimated models can be used as a decision oracle that informs the RL learning process and regulates RL policy updates cheng2019control . MPC’s regulatory capabilities are leveraged in the recently proposed RL-AR (Reinforcement Learning with Adaptive Regularization) tianreinforcement , an algorithm for safe training and deployment of RL that features convergence to the optimal RL policy. RL-AR comprises two parallel agents and an adaptive focus module: (i) The safety regularizer agent follows a deterministic policy $\pi_{\mathrm{reg}}:\mathcal{S}\rightarrow\mathcal{A}$ proposed by an MPC; (ii) The off-policy RL agent is an adaptive agent with $\pi_{\mathrm{rl}}:\mathcal{S}\rightarrow\mathcal{M}(\mathcal{A})$ that can learn from an acting policy that is different from $\pi_{\mathrm{rl}}$; (iii) The focus module learns a state-dependent weight $\beta:\mathcal{S}\rightarrow[0,1]$ for combining the deterministic $a_{\mathrm{reg}}(s)=\pi_{\mathrm{reg}}(s)$ and the stochastic $a_{\mathrm{rl}}(s)\sim\pi_{\mathrm{rl}}(s)$ actions proposed by the two agents. The safety regularizer component has a static built-in estimated environment model $\tilde{f}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ (that is different from the actual environment), while the off-policy RL agent and focus module are dynamically updated using observed interactions in the actual environment.
The RL-AR workflow is as follows: (i) $\pi_{\mathrm{reg}}(s)$ generates $a_{\mathrm{reg}}(s)$, which hard-codes safety constraints in the optimization problem over a period forecasted by $\tilde{f}$. The forecasting ability anticipates and prevents running into unsafe states for the critical system; (ii) $\pi_{\mathrm{rl}}(s)$ generates $a_{\mathrm{rl}}(s)$ to allow stochastic exploration and adaptation to the actual environment; (iii) $\beta(s)$ applies the following policy combination:
$\begin{split}a_{\beta}(s)=\beta(s)a_{\mathrm{reg}}(s)+(1-\beta(s))a_{\mathrm{rl}}(s).\end{split}$ (31)
The state-dependent focus weight $\beta(s)$ is initialized to $\beta(s)\geq 1-\epsilon$, $\forall s\in\mathcal{S}$, hence prioritizing the safe $\pi_{\mathrm{reg}}$ before $\pi_{\mathrm{rl}}$ learns a viable policy. As more interactions are observed for a state $s$ and the expected return of $\pi_{\mathrm{rl}}(s)$ improves, $\beta(s)$ gradually shifts the focus from the initially suboptimal $\pi_{\mathrm{reg}}(s)$ to $\pi_{\mathrm{rl}}(s)$. To update $\beta(s)$, the following objective is used:
$\begin{split}\beta^{\prime}(s)=\operatorname*{argmax}_{\beta\in[0,1]}\mathbb{E}\left[Q^{\pi_{\beta}}(s,\beta a_{\mathrm{reg}}(s)+(1-\beta)a_{\mathrm{rl}}(s))\right].\end{split}$ (32)
In tianreinforcement , it is analytically shown that if the reward function $R$ and transition probability $P$ of the MDP are Lipschitz continuous, then when a state $s$ has not been sufficiently explored and its corresponding $\beta(s)$ updates are limited, the sub-optimality of the RL policy $\pi_{\mathrm{rl}}$ has a limited impact on the expected return of the combined policy $\pi_{\beta}$, which is the actual acting policy. Meanwhile, as learning progresses and $\beta(s)$ approaches 0, RL-AR allows unbiased convergence to the optimal policy. RL-AR was shown effective in many safe RL applications ranging from medical to chemical systems, with a significant tolerance of estimated model deviation from the actual environment (see Section 4 in tianreinforcement ).
3.5 Learning State Dynamics
In the current integration of MPC with RL, the predominant framework in the control community is using discretized differential equations to describe the state evolution and thus hard physics defining boundaries of safety and other such considerations. Safety is guaranteed with the presence and enforcement, in a dedicated optimization procedure, of deterministic functional constraints of a single uniquely defined discretized dynamics, with the learning performed to inform the reward to objective layer.
However, consider the possibility that the physical dynamics have not been completely identified and quantified. Yet we stay a distance from the completely agnostic framework of NN representations, but instead consider that the equations have multiple subcomponent terms, some known exactly and some partially, and thus must be learned from data.
The broad framework can be referred to as Grey Box modeling. The purpose therein is to use any and all a priori physics and first principles based knowledge to formulate the dynamics of the process of interest. Meanwhile, the remaining unknown system dynamics is empirically modeled with data.
Indeed, all of the component functions in the MPC, including the dynamics $f$ and the constraints $c$, can be subject to learning, training and inference. The use of physics informed biases in the parametrized neural model for $f$ for instance could be used to enhance learning the precise differential system of equations.
As a recent example of NN-informed Grey-box identification, consider dong2022optimization ; dong2024descent . Broadly speaking, a powerful and effective approach is to consider a semilinear system,
$\mathcal{L}(u)+\mathcal{N}(u)=f,\,\nabla\cdot A\nabla u+\Phi(u;\theta)=f$
that is, $\mathcal{L}$ is a linear operator known to precision, and $\mathcal{N}$ is nonlinear and must be learned to, in general, produce a neural network model $\Phi(\cdot;\theta)$. For computing the solution of the MPC, the full estimated model is used to define the dynamics.
Robotics presents a natural application of hybrid RL-MPC online decision tools. Robotics has been a motivating application of RL as due to its enormous potential utility across a variety of domains. See for example the early survey kormushev2013reinforcement . Given their increasing role in automated manufacturer, home care, and other domains with vulnerable people or material, safety assurances become a priority for trusted stakeholder uptake. For a recent summary of work enforcing safety in robotics applications, see, e.g. brunke2022safe .
As a field of control, robotics presents a number of challenges as far as integrating the mechanical operative features in a manner that is accurate in modeling while also efficient with data. Entire fields of Mechatronics and Cybernetics include deep investigations as to the efficient and reliable interplay of physical mechanics, electronics, control and computing required for robot operation. One notable feature that can assist the sample complexity and degrees of freedom searched by solution methods is the presence of digitized movement choices, that is rotation along a finite set of distinct axes in freespace. This suggests the computational potential of scientific deep learning. Included is the use of geometric biases, e.g. the rotational movements of a robot often has a cyclic group structure which can be explicitly enforced in the modeling of the problem. The use of geometric learning and numerical methods can well be integrated in a control formulation, e.g. duruisseaux2023lie
4 Planning
4.1 Scheduling - Background
Scheduling is a fundamental problem in AI that involves allocating resources, time, and tasks to achieve specific objectives under given constraints dick2022scheduling . Formally, a scheduling problem consists of a set of tasks $T=\{t_{1},t_{2},\dots,t_{n}\}$, where each task $t_{i}$ has associated properties such as its duration $d(t_{i})$, start time $s(t_{i})$, and finish time $f(t_{i})=s(t_{i})+d(t_{i})$. Additionally, there is a set of resources (processors) $R=\{r_{1},r_{2},\dots,r_{m}\}$, where each resource $r_{j}$ may have capacity constraints $c(r_{j})$, limiting the number of tasks that can use $r_{j}$ simultaneously.
Constraints also play a critical role in scheduling problems. Precedence constraints ensure that certain tasks cannot start until others have finished, formalized as $s(t_{j})\geq f(t_{i})$, where $t_{i}$ must complete before $t_{j}$ begins. Resource constraints ensure that, at any time $t$, the number of tasks using a resource $r_{j}$ does not exceed its capacity $c(r_{j})$. Temporal constraints, such as lower bounds on start times, impose additional limits, for instance, $s(t_{i})\geq\tau_{i}$ for a given $\tau_{i}$.
The objective of a scheduling problem is to optimize a specific criterion (objective function). Common objectives include minimizing the makespan, defined as $\min\max_{t_{i}\in T}f(t_{i})$, which represents the total time required to complete all tasks dick2022scheduling . Other objectives include minimizing resource usage by reducing total cost or utilization and maximizing throughput, which aims to complete the maximum number of tasks within a given time limit.
Scheduling problems can be generally classified based on their structural properties. In single-machine scheduling, tasks are executed on a single resource, often with the goal of minimizing makespan or “lateness”. Parallel-machine scheduling extends this to multiple identical or unrelated resources baker2018principles . Job shop scheduling introduces a more complex structure, where tasks are grouped into jobs, each requiring a specific sequence of operations on different machines baker2018principles . Flow shop scheduling simplifies this by requiring that all jobs follow the same sequence of machines. More generally, constraint-based scheduling allows for arbitrary combinations of time, precedence, and resource constraints baptiste2001constraint .
Solving scheduling problems in AI is generally based on techniques from constraint satisfaction programming (CSP), linear programming, heuristic search, and metaheuristic optimization (e.g. genetic algorithms), which are employed to tackle the underlying combinatorial complexity dick2022scheduling . Applications of scheduling in AI then span various domains. In manufacturing, scheduling generally optimizes machine and resource allocation to maximize production efficiency. In industrial robotics, multi-robot systems rely on scheduling to allocate tasks while respecting time and resource limits. Staffing problems, such as in healthcare systems, are often framed as scheduling to allocate medical staff, equipment, and facilities, e.g. to maximize patient throughput and minimize wait times ernst2004staff .
4.2 Classical Planning - Background
Planning is another established framework for sequential decision making with a long tradition in AI ghallab2004automated . Planning shares core similarities with MPC and RL in their common objective of finding optimal action sequences. Like MPC, planning operates under the assumption of a known environment model, enabling proactive decision-making. Similar to most RL, planning focuses on discrete state and action spaces, facilitating the application of search algorithms. However, planning distinguishes itself through its goal-centric approach and the use of relational representations of states and actions, a key difference from the numerical representations typically used in MPC and the enumerated representations common in (tabular) RL. This relational approach allows planning to handle more complex, abstract relationships and generalize over different objects and environments more effectively.
4.2.1 Relational Logic
First-order (predicate) logic is the cornerstone of all formal knowledge representations and reasoning in AI, providing the formal language for expressing facts, rules, and relationships between concepts huth2004logic .
Syntactically, in first-order logic (FOL), we define a vocabulary (signature) $\Sigma$ consisting of: a set of constant symbols $\mathcal{C}$ (e.g., “Room1”,“Patient2”); a set of variable symbols $\mathcal{V}$ (e.g., position “$x$”, distance“$d$”); a set of function symbols $\mathcal{F}$ (e.g., “$\text{neighborOf}(x)$”), each with an associated arity $n\in\mathbb{N}$ denoted by $f/n$ where $f\in\mathcal{F}$; and a set of predicate symbols $\mathcal{P}$ (e.g., “$\text{Asleep}(x)$” or “$\text{Nursing}(x,y)$”), also with associated arities $P/n$ where $P\in\mathcal{P}$. The set of terms $\mathcal{T}(\Sigma)$ over $\Sigma$ is then defined recursively: $\mathcal{C}\subset\mathcal{T}(\Sigma)$, $\mathcal{V}\subset\mathcal{T}(\Sigma)$; and if $f/n\in\mathcal{F}$ and $t_{1},\dots,t_{n}\in\mathcal{T}(\Sigma)$, then $f(t_{1},\dots,t_{n})\in\mathcal{T}(\Sigma)$ (e.g., “nurseOf(Patient1)”). The set of formulae $\mathcal{F}(\Sigma)$ over $\Sigma$ is then also defined recursively as follows. If $P/n\in\mathcal{P}$ and $t_{1},\dots,t_{n}\in\mathcal{T}(\Sigma)$, then $P(t_{1},\dots,t_{n})\in\mathcal{F}(\Sigma)$ is an atomic formula; if $\phi,\psi\in\mathcal{F}(\Sigma)$, then $\neg\phi,(\phi\land\psi),(\phi\lor\psi),(\phi\rightarrow\psi),(\phi\leftrightarrow\psi)\in\mathcal{F}(\Sigma)$, and if $\phi\in\mathcal{F}(\Sigma)$ and $x\in\mathcal{V}$, then $\forall x\phi$ and $\exists x\phi\in\mathcal{F}(\Sigma)$ (e.g., “$\forall x(\text{Man}(x)\rightarrow\text{Mortal}(x))$”).
Semantics in logic is then defined by the means of interpretations huth2004logic . An interpretation $\mathcal{I}=(D,I)$ consists of a non-empty domain $D$ and an interpretation function $I$: for each $c\in\mathcal{C}$, $I(c)\in D$; for each $f/n\in\mathcal{F}$, $I(f):D^{n}\rightarrow D$; and for each $P/n\in\mathcal{P}$, $I(P):D^{n}\rightarrow\{\text{True},\text{False}\}$. A variable assignment (substitution) $v:\mathcal{V}\rightarrow D$ then maps variables to elements in $D$. The interpretation of a term $t$ under $\mathcal{I}$ and $v$ is denoted $t^{\mathcal{I},v}$. Truth of a formula $\phi$ is denoted $\mathcal{I},v\models\phi$. $\mathcal{I}$ satisfies $\phi$, denoted $\mathcal{I}\models\phi$, if $\mathcal{I},v\models\phi$ for all $v$.
The relational logic is then an important function-free ($\mathcal{F}=\emptyset$) subset of FOL that underlies the representation formalism of classical planning.
4.2.2 Planning
The problem of planning is normally formalized in the language of predicate logic huth2004logic . The world is represented by a set of predicates $\mathcal{P}$, where each $p\in\mathcal{P}$ has arity $\text{arity}(p)$. A state $s\in\mathcal{S}$ is then a truth assignment $s:\mathcal{P}\to\{\text{True},\text{False}\}$. For a ground predicate $p(c_{1},\dots,c_{\text{arity}(p)})$ over the respective constants $c_{i}$, $s(p(c_{1},\dots,c_{\text{arity}(p)}))$ denotes its truth value.
Actions (controls) are operators $a\in\mathcal{A}$ with precondition $\text{pre}(a)$ and effect $\text{eff}(a)=(\text{eff}^{+}(a),\text{eff}^{-}(a))$, representing added and deleted predicates. Action $a$ is applicable in $s$ if $s\models\text{pre}(a)$. Applying $a$ to $s$ results in a new state $s^{\prime}={\psi}(a,s)$:
$s^{\prime}(p)=\begin{cases}\text{True}&\text{if }p\in\text{eff}^{+}(a)\\ \text{False}&\text{if }p\in\text{eff}^{-}(a)\\ s(p)&\text{otherwise}\end{cases}$ (33)
A planning problem is defined by an initial state $s_{0}$ and a goal condition $\phi_{G}$, represented generally by a logical formula. A plan $\mathbf{a}=(a_{0},\dots,a_{n-1})\in\mathcal{A}^{n}$ achieves the goal if applying it to $s_{0}$ results in a state $s_{n}$ satisfying $\phi_{G}$:
$s_{n}\models\phi_{G},\quad\text{where }s_{t+1}=\psi(a_{t},s_{t}),\quad t=0,\dots,n-1.$ (34)
Planning algorithms can be generally categorized as either satisficing or (less commonly) optimal, depending on whether they guarantee to achieve the goal or also finding the shortest plan. Somewhat similar to MPC (Eq. 7), optimal planning can be formulated as aiming to minimize a cost:
$\mathbf{a}^{*}=\arg\min_{\mathbf{a}}C(\mathbf{a}),\quad C(\mathbf{a})=\sum_{t=0}^{n-1}C(s_{t},a_{t})+C_{T}(s_{n}),$ (35)
subject to the state transitions defined by Eq. (33) and the goal satisfaction (34). The state transition function $\psi$ (Eq. (33)) can be thought of analogously to MPC (Eq. 7) and RL, but operates on relational structures instead of numerical vectors. Similarly, the reward function $R$ or cost $C$ is analogous to the cost function in optimal planning, but planning explicitly focuses on achieving the goal (terminal) state, which is most commonly the only reward. Connecting these fields further, model-based RL moerland2023model often utilizes planning by learning a model and then using it for search, bridging the gap between data-driven learning and model-based control.
Conceptually, classical planning could be viewed as a special case of MPC where (i) the transition dynamics ($\mathcal{T}(s,a)$) are discrete and deterministic, (ii) the cost function prioritizes goal achievement (typically 00 if $x\models\mathcal{G}$ otherwise $\infty$), and (iii) the optimization horizon spans until the goal is reached.
4.3 Planning extensions (Multicriteria and Online Planning)
As the classical planning lacks the expressiveness needed for real-world dynamic systems, there are also extensions such as numeric, temporal, and stochastic planning that bring it a bit closer to the dynamic, real-time nature of MPC. Numeric planning incorporates continuous state variables and numeric fluents, partially enabling modeling dynamic systems akin to MPC’s difference equations. Temporal planning introduces explicit reasoning about time, including action durations and deadlines. Stochastic planning then addresses uncertainty by modeling probabilistic (or worst-case) transitions.
Also, classical PDDL (Planning Domain Definition Language) lavalle2006planning planning typically involves finding a sequence of actions $\pi=\langle a_{1},a_{2},\dots,a_{T}\rangle$ that transitions the system from an initial state $s_{0}$ to a goal state $s_{n}\models\phi_{G}$. Extensions following the numeric and optimal planning then led to introducing plan metrics $F(\pi)$, functions that evaluate the quality of a plan based on numeric fluents $\{f_{1},f_{2},\dots,f_{k}\}$ in the domain (e.g., fuel, distance, etc.). These metrics allow differentiating between plans that achieve the same goal $G$ but with varying numeric performance. For example, a plan $\pi$ is optimal if $\pi^{*}=\arg\min_{\pi}F(\pi)$, where $F(\pi)$ might minimize action cost, time, or resource usage. The metric $F$ can combine the multiple criteria through functions over fluents, such as $F(\pi)=\sum_{i=1}^{k}w_{i}F_{i}(\pi)$, yielding context-dependent optimal plans. This approach, more reminiscent of the function optimization in MPC and RL, is supported in extended PDDL versions, however, remains rarely supported by actual planners.
Further, classical PDDL planning operates under the assumption of a static, fully observable environment, generating a complete plan $\pi=\langle a_{1},a_{2},\dots,a_{T}\rangle$ offline to transition from an initial state $s_{0}$ to a goal state $s_{g}$, where $s_{g}\models G$ for a fixed goal condition $G$. However, real-world environments often exhibit and unpredictable behavior, which can be addressed by interleaving planning and execution to achieve a form of “online planning.” This can be modeled through iterative replanning, where the system updates the problem state $s_{t}$ to reflect the current environment and recomputes the optimal plan $\pi^{*}=\arg\min_{\pi}F(\pi\mid s_{t})$ from the current $s_{t}$. A possible hybrid approach is then to combine PDDL with inherently dynamic frameworks like the MPC or RL, where PDDL handles high-level planning, and the dynamic framework manages the uncertainties by optimizing policies $\pi^{*}$ over the underlying state transitions $P(s_{t+1}\mid s_{t},a_{t})$ in real-time.
4.4 Planning (and RL) with (Graph and) Relational Neural Networks
Classical planning has been extended in recent years to attempt to leverage the computational and representational power of neural modeling and training. As in MPC, the techniques of RL can be used to learn both the functions defining the classical planning problem, the transition and cost, through data, or even be used to compute an optimal policy. Before continuing to describe some of the main methodological tools available, we also indicate another contemporary AI tool component: the use of structured data representations. By using Graphs, and more generally structured relations, we can endow an inductive bias into a wide variety of use cases that facilitates computation as well as interpretability.
Recent advancements in structured deep learning models, particularly Graph Neural Networks (GNNs) zang2023graphmp , have proven effective for classical planning tasks, which can often be naturally encoded in the form of various graphs. Given a graph $G=(V,E)$, where $V$ represents entities, such as domain objects, and $E$ represents relationships between them, a GNN learns the embeddings of nodes $\mathbf{v}_{i}^{(K)}$ by iteratively passing messages between nodes. The message-passing update rule is generally given by:
$\mathbf{v}_{i}^{(k+1)}=\phi\left(\mathbf{v}_{i}^{(k)},\bigoplus_{j\in\mathcal{N}(i)}\psi(\mathbf{v}_{j}^{(k)},\mathbf{e}_{ij})\right),$ (36)
where $\mathcal{N}(i)$ denotes the neighbors of node $i$, $\mathbf{e}_{ij}$ is the, possible feature-associated, edge between nodes $i$ and $j$, $\psi$ is the message function, and $\phi$ updates the node embedding. After $K$ iterations, the learned node embeddings $\mathbf{v}_{i}^{(K)}$ are aggregated and used to predict heuristics $h(s)$, which are then used in search algorithms. The most prominent in planning, i.e. searching for a sequence of actions that leads from an initial state $s_{0}$ to a goal state $s_{g}$, is the classic A* search that maintains a priority queue of states $s$, ordered by the evaluation function $f(s)$, which is defined as $f(s)=g(s)+h(s)$, where $g(s)$ is the known cost to reach state $s$ from the initial state $s_{0}$, and $h(s)$ is the heuristic estimate of the cost to reach the goal state $s_{g}$. At each step, A* selects the state with the lowest $f(s)$, expands its neighbors, and updates their $f$-values accordingly. With enough data samples, the GNN can often learn a heuristic $h(s)$ based on the graph structure of the planning problem, improving the A* search efficiency by providing more accurate cost-to-go estimates, leading to fewer expanded nodes and faster planning. Structured neural networks can also be applied to learn the integration of scheduling and planning ma2020online
Relational NNs
However, standard GNNs can not integrate domain’s background knowledge (BK) directly into the learned model heuristics. To address this, one may use more expressive “neuro-symbolic” frameworks based on the language of relational logic (Sec. 4.2.1). One such prominent example framework are Lifted Relational Neural Networks (LRNNs) sourek2018lifted , which extend GNNs by incorporating relational rules that guide the learning process. Similarly to RL, a planning policy $\pi(s)$ can be a (learned) model that selects the (optimal) action $a$ given a state $s$, i.e.,
$\pi(s)=\arg\max_{a\in A}C(a\mid s),$ (37)
where $A$ is the set of all possible actions, and $C(a\mid s)$ is some (learned) scoring function. With LRNNs, however, domain knowledge can also be encoded, e.g. in the form of a prior nondeterministic policy $\sigma(s)$ that returns merely a subset of feasible actions $\sigma(s)\subseteq A$. Similarly to the T-S fuzzy control formalism (Sec. 5), this policy can be defined by rules of the form:
$\text{condition}(s)\to\text{action}(a),$ (38)
where $\text{condition}(s)$ is a relational logic formula over the state $s$, and $\text{action}(a)$ specifies the feasible actions according to the domain knowledge capturing some desired domain constraints, such as safety. This nondeterministic policy $\sigma(s)$ then effectively restricts the learning hypothesis space of the model $\pi(s)$. For instance, in the popular Blocksworld domain lavalle2006planning , $\sigma(s)$ may specify that misplaced blocks are unstacked onto the table before restacking them in the correct order, which ensures a satisficing but potentially suboptimal policy. The policy $\pi(s)$ can then be further refined by optimizing the scoring function $F(a\mid s)$ that ranks actions in $\sigma(s)$ based on their likelihood of improving the given plan quality metrics, such as minimizing cost. This allows the model to focus on optimizing the solution quality while adhering to the given BK constraints, instead of solving the planning problem from scratch in potentially unsafe ways.
In general, integrating logical (relational) background knowledge can be used for domain-guided planning, where learned policies $\pi(s)$ are shaped by both training data and expert-provided domain knowledge. This offers advantages over existing, e.g. GNN-based, learning-to-plan methods by ensuring that the learned policies are both satisficing and aligned with real-world constraints.
Reinforcement Learning with Logic and Relational Policies
Related ideas have also been explored in the reinforcement learning setting with other neuro-symbolic frameworks, such as $\delta$ILP evans2018learning , which aims to relax the problem of Inductive Logic Programming (ILP), or rule learning, into a differentiable setting via (constrained) rule templates where each rule is associated with a learnable weight, similar to the LRNNs, and trained with gradient descent. Building on that, frameworks such as jiang2019neural integrate $\delta$ILP with RL to enhance interpretability and generalization through the logical formalism. The RL problem is modeled as an MDP with logic components $(\mathcal{M},p_{S},p_{A})$, where $\mathcal{M}=(\mathcal{S},\mathcal{A},T,R)$ defines the state space, action space, transitions, and rewards. A state encoder $p_{S}:\mathcal{S}\to 2^{\mathcal{G}}$ maps states to sets of logical atoms, and an action decoder $p_{A}:[0,1]^{|\mathcal{D}|}\to[0,1]^{|\mathcal{A}|}$ translates the logic valuations into action probabilities. Using $\delta$ILP, the policies are represented as valuation vectors $e\in[0,1]^{|\mathcal{G}|}$, where each element indicates the confidence of a logical atom being true. Inference then occurs via a recurrent neural model $f_{\theta}$, which iteratively updates the valuations through time. Logical rules are encoded as weighted first-order clauses of the form
$\alpha\leftarrow\alpha_{1},\ldots,\alpha_{n},$
where $\alpha$ is the head atom, and $\alpha_{1},\ldots,\alpha_{n}$ are body atoms, again similarly to LRNN or the T-S fuzzy systems. Trainable clause weights, constrained by a softmax function, then ensure differentiability, making the policy
$\pi(s)=p_{A}(f_{\theta}(p_{S}(s))),$
combining logic reasoning with standard RL optimization using policy gradients to maximize the expected returns.
4.5 Learning for Planning
The previous section described how to integrate the crisp structure planning provides as far as representations of policies into an otherwise data-driven scheme for performing Reinforcement Learning.
Here, consider now the converse, that one has a planning problem, and because of unknown functions, learning must be performed. We can see the resemblance to Dynamic Programming again as with MPC. Rather, planning shared the discrete action and state space of classical DP, unlike MPC. However, although an objective function can be present in both, the presence and moreover priority of the goal state is particular to planning.
Consider again the components of the planning problem functions. Recall the equations associated with the planning problem:
$s_{n}\models\phi_{G},\quad\text{where }s_{t+1}=\psi(a_{t},s_{t}),\quad t=0,\dots,n-1.$
$\mathbf{a}^{*}=\arg\min_{\mathbf{a}}C(\mathbf{a}),\quad C(\mathbf{a})=\sum_{t=0}^{n-1}C(s_{t},a_{t})+C_{T}(s_{n},n),$
Now, consider that, unlike in classical planning, the functions $\psi$ and $F$ are not known, but must be learned. This can be done by observing the sequences of real states $\{\hat{s}_{1},\hat{s}_{2},\cdots,\hat{s}_{n}\}$ as well as obtaining rewards $R_{t}$ during the process. For learning, parametrization of the problem functions can be performed,
$\psi(a_{t},s_{t}):=\psi(\theta^{\psi},a_{t},s_{t}),\,F(a_{t},s_{t}):=F(\theta^{F},a_{t},s_{t}),\,F_{T}(s_{n},n):=F_{T}(\theta^{T},s_{n},n)$
and these functions can be learned directly, which would correspond to a model based approach of reinforcement learning.
Alternatively, a purely model-free approach can solve for the solution of the planning problem without learning the components of the function. In the standard way, an action value function $V(s_{t},a_{t})$ can be directly learned, or simply a policy gradient approach used, to compute and define a policy
$\pi(s)=p_{A}(f_{\theta}(p_{S}(s))),$
as described above. In this case, however, the original structure of the planning problem is hardly maintained, i.e., this approach is barely distinguishable from standard RL.
Later in this work, we do consider uncertain realizations of the dynamics and rewards in the context of care planning. However, we will use a lower level state representation, describing detailed physical dynamics, and run by an MPC-based solver, to assist in, ultimately more accurate and robust, data-driven descriptions of the planning operation and its solution.
5 Fuzzy Sets and Fuzzy Control - Background
Fuzzy sets and fuzzy control provide a mathematical framework for dealing with imprecision and vagueness. Traditional set theory, based on binary membership where an element either belongs to a set or not, is often insufficient for capturing the vagueness inherent in many real-world systems. Fuzzy set theory, introduced by zadeh1965fuzzy , extends the classical sets by allowing elements to have varying degrees of membership, thereby modeling the uncertainty and imprecision.
5.0.1 Fuzzy Sets
A fuzzy set $\tilde{A}$ in some universal set of discourse $X$ is defined by its membership function $\mu_{\tilde{A}}:X\to[0,1]$, where $\mu_{\tilde{A}}(x)$ indicates the degree to which an element $x\in X$ belongs to $\tilde{A}$. Formally, a fuzzy set can be represented as:
$\tilde{A}=\{(x,\mu_{\tilde{A}}(x))\mid x\in X\},$ (39)
where $\mu_{\tilde{A}}(x)=1$ indicates full membership, $\mu_{\tilde{A}}(x)=0$ indicates non-membership, and $10<\mu_{\tilde{A}}(x)<$ 1 represents a partial membership. Common examples of membership functions include triangular, trapezoidal, Gaussian, and sigmoid functions bai2006fundamentals .
Basic set operations on these fuzzy sets are then commonly defined as
  • Union: $\mu_{\tilde{A}\cup\tilde{B}}(x)=\max(\mu_{\tilde{A}}(x),\mu_{\tilde{B}}(x))$
  • Intersection: $\mu_{\tilde{A}\cap\tilde{B}}(x)=\min(\mu_{\tilde{A}}(x),\mu_{\tilde{B}}(x))$
  • Complement: $\mu_{\tilde{A}^{c}}(x)=1-\mu_{\tilde{A}}(x)$
In practice, fuzzy sets are then used to model “linguistic variables”, where terms such as “low, medium, or high” are represented by fuzzy sets with appropriately defined membership functions. For example, in a temperature control system, the variable “temperature” could be modeled using fuzzy sets such as Low, Medium, and High.
5.0.2 Fuzzy Logic
Fuzzy logic then extends classical Boolean logic by incorporating degrees of truth. In fuzzy logic, the truth value of a statement is again not restricted to binary values (00 or $1$) but can take any value in the interval $[0,1]$. To that aim, fuzzy logic operators extend classical logical operators to operate on the fuzzy sets bai2006fundamentals .
These fuzzy logic operators are then typically defined using t-norms $T:[0,1]^{2}\rightarrow[0,1]$ (for conjunction, satisfying commutativity, associativity, monotonicity, and having 1 as identity) and t-conorms $S:[0,1]^{2}\rightarrow[0,1]$ (for disjunction, with analogous properties). Common choices include: minimum t-norm ($T(a,b)=\min(a,b)$), product t-norm ($T(a,b)=a\cdot b$), Łukasiewicz t-norm ($T(a,b)=\max(0,a+b-1)$), maximum t-conorm ($S(a,b)=\max(a,b)$), probabilistic sum t-conorm ($S(a,b)=a+b-a\cdot b$), and Łukasiewicz t-conorm ($S(a,b)=\min(1,a+b)$). Negation is then commonly defined as $\neg a=1-a$, although there are different choices as well.
5.0.3 Fuzzy control
Fuzzy control utilizes fuzzy set theory and fuzzy logic to design control systems passino1998fuzzy . A fuzzy controller typically consists of four main components. First, Fuzzification involves converting crisp input values into fuzzy sets using the membership functions. Second, Inference applies fuzzy rules to evaluate outputs based on fuzzy inputs. Third, Aggregation combines the results from all applicable rules to form a single fuzzy output set and, fourth, Defuzzification converts the fuzzy output set into a crisp output value.
The fuzzy rules are typically expressed in the form of “IF-THEN” statements, where the antecedent (IF part) describes the input conditions using fuzzy sets, and the consequent (THEN part) determines the output action.
Mamdani system
The Mamdani fuzzy inference system mamdani1975experiment is one of the earliest and most widely used fuzzy control methodologies. It is based on the idea of representing both antecedents and consequents as fuzzy sets. A typical rule in the Mamdani model is expressed as:
IF $\text{IF }x_{1}\text{ is }\tilde{A}_{1}\text{ AND }x_{2}\text{ is }\tilde{A}_{2}\text{ THEN }y\text{ is }\tilde{B},$ (40)
where $x_{1},x_{2}$ are input variables, $\tilde{A}_{1},\tilde{A}_{2}$ are fuzzy sets for the antecedents, and $\tilde{B}$ is the fuzzy set describing the consequent.
In Mamdani systems, the inference process proceeds as follows. For each rule, the degree of activation is computed by evaluating the membership functions of the antecedents. Commonly, the min operator is used to compute the degree of fulfillment for the conjunction (t-norm) of antecedents (joined by AND), i.e.,
$\mu_{\tilde{R}_{i}}(x_{1},x_{2})=\min(\mu_{\tilde{A}_{i1}}(x_{1}),\mu_{\tilde{A}_{i2}}(x_{2})).$ (41)
The output fuzzy set $\tilde{B}_{i}$ of each rule is then scaled by the degree of activation $\mu_{\tilde{R}_{i}}$, e.g. using again the min operator:
$\mu_{\tilde{B}_{i}}(y)=\min(\mu_{\tilde{B}}(y),\mu_{\tilde{R}_{i}}).$ (42)
The scaled output fuzzy sets $\tilde{B}_{i}$ from all rules are then combined, typically using the max operator:
$\mu_{\tilde{B}_{\text{agg}}}(y)=\max_{i}\mu_{\tilde{B}_{i}}(y).$ (43)
Finally, the aggregated fuzzy output set $\tilde{B}_{\text{agg}}$ is converted into a crisp value through defuzzification. A common method is the centroid method, where the crisp output $y^{*}$ is computed as:
[…]
The Mamdani approach is particularly intuitive and well-suited for human-interpretable control systems, where the rules are derived from expert knowledge.
Takagi-Sugeno system
The Takagi-Sugeno (T-S) fuzzy model takagi1985fuzzy is the other prominent formalism that takes a different approach by employing crisp mathematical functions in the consequents of the rules. A typical T-S rule is expressed as:
IF $\text{IF }x_{1}\text{ is }\tilde{A}_{1}\text{ AND }x_{2}\text{ is }\tilde{A}_{2}\text{ THEN }y=g(x_{1},x_{2}),$ (45)
where $g(x_{1},x_{2})$ is typically a linear function:
$g(x_{1},x_{2})=a_{1}x_{1}+a_{2}x_{2}+b,$ (46)
with $a_{1},a_{2},b\in\mathbb{R}$ as parameters.
In the T-S model, the inference process begins by computing the degree of activation $\mu_{\tilde{R}_{i}}(x_{1},x_{2})$ for each rule, as in the Mamdani system. The weighted output for each rule is then calculated based on its activation degree:
$y_{i}=\mu_{\tilde{R}_{i}}(x_{1},x_{2})\cdot g(x_{1},x_{2}).$ (47)
The outputs of all rules are combined using a weighted average:
$y^{*}=\frac{\sum_{i=1}^{n}\mu_{\tilde{R}_{i}}(x_{1},x_{2})\cdot g_{i}(x_{1},x_{2})}{\sum_{i=1}^{n}\mu_{\tilde{R}_{i}}(x_{1},x_{2})}.$ (48)
The T-S model is computationally efficient and well-suited for control problems where a mathematical model of the system dynamics is available.
6 Integrating the Symbolic and Control Layers
Consider now the robotic aide described in the Introduction of the manuscript. Here we describe an integration of the low level control operation with the higher level conceptual discrete planning and scheduling decision making.
Figure 1 presents the high level picture of the operation. Over a long, slow time scale, a Scheduler decides upon a sequence of tasks. When the time comes for each task, a planning problem corresponding to that task is formulated and solved. In the course of executing the solution of the planning problem, each action taken generates a continuous control problem that is solved at a fast time scale that yields the true reward and next state, which may or may not be the one predicted by the planner. Models of rewards and transitions are learned, with information flowing across layers in a continuous, sequential manner.
Figure 1: A pictorial model of integrating the Symbolic Planning and Control Layers in Operation. The scheduler decides upon the sequence of Tasks the robot is set to perform. Each Task defines a Planning Problem. The solution to the problem is computed, and then at each time step, for a given state an action is taken which then defines an MPC problem. The MPC problem is solved in real time, and the final (continuous) state is mapped to the set of (discrete) states in the planning ontology. Discrepancies in the form of intermediate rewards and unexpected state transitions can change the planning problem in real-time, necessitating a re-computation of the planning solution.
6.1 Hierarchical Planning-MPC Formulation
Consider the original formulation of the planning problem, wherein a set of actions $a(t)\in\mathcal{A}$ with $\left|\mathcal{A}\right|=A$ must be chosen in order for the state $s_{t}\in\mathcal{S}$ with $\left|\mathcal{S}\right|=S$ to proceed in order to satisfy a goal state $x(T)\in\mathcal{G}$ while minimizing some cost. The process is defined as a stochastic planning problem, that is, now the cost functions and state transitions include a random variable. For a contemporary example, see flores2020stochastic . While the challenging field is growing, it is significantly less mature than classical planning itself. Moreover, we will assume limited probabilistic knowledge a priori as to the distributions of these functions, and instead focus on a data driven approach for learning these functions. Specifically, we will use parametric function approximations for defining the expectations of the costs.
$\begin{array}[]{l}C(\mathbf{s},\mathbf{a})=\sum_{t=0}^{n-1}C(s_{t},a_{t})+C_{T}(s_{n})+\rho(n),\\ C(s,a)=\mathbb{E}_{\xi}[c(s,a,\xi(s,a))\\ C_{T}(s)=\mathbb{E}_{\xi}[c_{T}(s_{n},\xi(s_{n},a_{n}))]\end{array}$ (49)
where $\rho(n)$ is monotonically increasing with respect to $n$ and is a time penalty of the number of steps until the goal state is reached. The cost, which will depend on some metric of the overall well-being of the human and agreeableness of the robotic service, is treated as stochastic. Note that the randomness itself depends on the state and action taken, modeling the potential variability of, e.g., emotional affect in response to a robot’s action across different circumstances.
Consider that the transitions are governed by the deterministic but unknown dynamics
$s_{t+1}=\psi(s_{t}(z_{t}),a_{t},z_{t},e_{t})$
where $z_{t}$ is a continuous state of the robot and $e_{t}$ is an exogenous set of quantities that affect the transition process. Observe that the complete continuous state $z_{t}$ completely identifies the discrete symbolic state $s_{t}$ and $e_{t}$ is unpredictable and only partially observed. Thus, the actual state transitions are unidentifiable, and the planning computation is performed with a surrogate. The fact that these plans are executed multiple defines in an online sequence set by a scheduler makes the problem tractable. That is, with a one off planning problem, one clearly cannot learn the terminal state cost. However, since each plan is expected to be repeated and a scheduler will define tasks into an indefinite future time period, these function approximations are expected to become increasingly more accurate over time.
This surrogate will be defined by function approximations, introducing parameters $\theta^{C}$ and $\theta^{T}$ for the cost functions and $\theta^{\psi}$ for the state transition. Specifically, the functional forms for the model the operator uses to compute the plan as:
$\begin{array}[]{l}\bar{C}(s_{t},a_{t})=\Phi^{C}(\theta^{C};s_{t},a_{t})\\ \bar{C}_{T}(s_{n})=\Phi^{T}(\theta^{T};s_{n})\\ \bar{s}_{t+1}=\bar{\psi}(s_{t},a_{t})=\max\limits_{s\in\mathcal{S}}p_{\psi}^{S}(\theta^{\psi};s_{t},a_{t})\end{array}$ (50)
that is, the cost function is some parametrized model with inputs corresponding to the action, previous and subsequent states, and intermediate process, and the state transition is predicted $\bar{s}_{t+1}$ by a probability density also depending on these quantities. While stochastic planning is a possible approach to handling uncertainty, we prefer to use the more comprehensive and reliable toolkit of classical planning algorithms and simply choose perform a softmax of the distribution of the predicted state.
We will denote the sequence $\{\hat{s}_{t},\hat{a}_{t}\}$ to be the physically realized action and state pair that the procedure executes and transitions through. In addition, at current time $\hat{t}$, the intermediate sequence of reward has been observed as $\{\hat{C}_{0},\hat{C}_{1},\cdots,\hat{C}_{\hat{t}-1}\}$. We consider that the same planning problem will be solved multiple times by the robot and so historical rewards and states are also available and have informed the models $\bar{C}$, $\bar{C}_{T}$ and $p_{\psi}^{S}$.
The lower level control process is defined as follows: in between each transition $t$ to $t+1$, a state-control process running at a faster time scale $\{z_{\tau},v_{\tau}\}$ with $\tau$ ranging from $t,t+h,t+2h,\cdots,t+1$ with $h\ll 1$. We remark that while the planning state and action $(s_{t},a_{t})$ are members of a discrete set, the lower level control operation $\{z_{\tau},v_{\tau}\}$ is defined by continuous vectors. Naturally, there is a correspondence between the continuous and discrete state, which arises in the overall operation loop when the next state $\hat{s}_{t+1}$ is inferred from the final state in the control problem $\hat{z}_{t+1}$, which we define later. As a simple example, $s_{t}$ can indicate which out of a finite set of rooms a robot is in, and $z_{\tau}$ the precise coordinates of its location.
We note that nonuniform and adaptive step-sizes, multiple shooting formulations, etc. can be appropriately defined with natural extensions of the formulation.
Below we proceed to detail the functional components of our procedure of integrating the planning, MPC, and learning aspects of the procedure.
6.2 Selecting the planning action
We consider real time closed loop planning, in which not only is a plan computed at the initial moment of operation, but for each step in the operating sequence, the plan is recomputed. This is necessary because of the model of the functions used for planning, that is $\Phi^{V}$ and $p^{S}_{\psi}$ are updated at each step to reflect discrepancy in the expected and real reward, that is $\bar{C}(s_{t},a_{t})$ compared to $\hat{C}_{t}$, as well as the transition model by $\bar{s}_{t+1}$, the predicted state, relative to $\hat{s}_{t+1}$, the actual realized state. For the rest of this section, consider a consider that the current time of real-time online operation is given by $\hat{t}$.
In this section we present several options with respect to how one computes a plan at each time $\hat{t}$. To this end we assume that at the current time $\hat{t}$ all of the model functions are current, i.e., have been updated with the most recent observations.
At this stage the operation has updated information that may affect the optimal plan that was computed at $\hat{t}-1$. This necessitates that the planner should then recompute the plan for the remaining time $\{\hat{a}_{\hat{t}},\bar{s}_{\hat{t}+1},\bar{a}_{\hat{t}+1},\bar{s}_{\hat{t}+2},\bar{a}_{\hat{s}+2},\cdots\bar{s}_{n}\}$. Here the quantities $\bar{s}$ and $\bar{a}$ indicated the predicted future states and the computed actions for these predicted states. The action chosen to be implemented once $\hat{t}\geq t$ is given as $\hat{a}_{t}$. For the choice of algorithm to perform this computation, there are several options which will be outlined below.
Actions as RL Policies
A classic RL approach would correspond to the choice $\hat{a}_{t}$ being computed in accordance with some action-value function $V(\theta^{V},a,s)$. This is meant to represent a cost-to-go, that is,
$V(\theta^{V},a_{t},s_{t})=\max\limits_{a\in\mathcal{A}}\left\{\begin{array}[]{ll}&-\bar{C}(s_{t},a)-\sum\limits_{t=\hat{t}+1}^{n-1}\bar{C}(\bar{s}_{t},\bar{a}_{t})-C_{T}(\bar{s}_{n})-\rho(n)\\ \text{such that }&\bar{s}_{n}\in\mathcal{G},\\ &\bar{s}_{t+1}=\bar{\psi}(\bar{s}_{t},\bar{a}_{t}),\,t>\hat{t}\end{array}\right\}$ (51)
Observe that learning the intermediate and terminal cost functions czn be used to define the action-value function, or $\theta^{V}$ can be used to approximate the model on the right, eschewing the need for $\theta^{C}$, etc.
The standard techniques here correspond to greedy, $\epsilon$-random greedy, and logit:
$\begin{array}[]{lr}\hat{a}_{t}=\hat{a}_{t}^{g}=\arg\max_{a}V(\theta^{V}_{t},a,s_{t})&\text{greedy}\\ \hat{a}_{t}=\hat{a}^{\epsilon}_{t}=\left\{\begin{array}[]{lr}\arg\max_{a}V(\theta^{V}_{t},a,s_{t})&\text{w.p.}(1-\epsilon)\\ a^{\prime}&\text{w.p.}\frac{\epsilon}{A},\,\,\forall a^{\prime}\in\mathcal{A}\end{array}\right.&\epsilon-\text{greedy}\\ \hat{a}_{t}=\hat{a}^{\sigma}_{t}=\hat{a}(\omega^{a}_{t})\sim_{a}e^{-V(\theta^{V}_{t},a,s_{t})}/Z&\text{logit}\end{array}$ (52)
Classical Planning on the Learned Model
Alternatively, some classical planning algorithm $\mathcal{C}$ can perform a selection of $\hat{a}_{t}$ based on purely logical representations, treating the learned models as fixed exogenously at the point of computation:
$\hat{a}^{c}_{t}=\mathcal{C}(\bar{\psi},\bar{C},\bar{C}_{T},\rho,s_{t}),$
where $\mathcal{C}$ may be a learned policy, that is a neural map from the parametrization of the models to the action sequence, or a search procedure, driven by the respective goal satisfaction and cost minimization. Various exact and heuristic search methods, such as the well known A* procedure, can be used to search for a policy.
This approach is most analogous to the MPC-RL integration described above, as far as the use of classical numerical procedures with firm theoretical guarantees being applied to compute the solution, while neural networks are used to parametrize and model the component objective, or value, functions.
6.3 Action-Triggered MPC
Consider a real time operation of planning with $\hat{t}$ denoting the current time. A plan $\zeta(s)$ that at time $\hat{t}$ with $\hat{t}>0$ will execute a particular action $\hat{a}_{t}{\hat{t}}$. This action triggers the solution and execution of an MPC problem. The action defines the problem functions, that is the dynamics, objective, acceptable continuous actions, and any constraints, and the MPC problem is solved in real-time closed loop fashion, delivering a control sequence. This control sequence transitions the continuous state until the terminal state, which is then mapped onto an appropriate discrete state. In addition, any additional reward-relevant information that appeared during the course of the control solution is passed to the planner to assist with updating a more faithful plan at the next discrete time instance.
Formally, the MPC final state is expected to yield, upon its mapping onto the discrete set $\mathcal{S}$, the predicted state that was associated with the choice of action $a_{\hat{t}}$. This is based on the model at time $\hat{t}$, $p^{S}_{\psi}(\theta^{\psi};s_{t},a_{t})$, this defines the predicted state $\bar{s}_{t+1}$ at the next time step. The details of how $\theta^{\psi}$ is learned and $\bar{s}_{t+1}$ is computed will be the topic of subsequent subsections.
We define the action of the realized $\hat{a}_{t}$ to be the formulation and approximate solution of a discretized optimal control problem. The form of the objective $G$ and $W$, dynamic state equation $f$ and constraints $h$ defined in the NLP below depend on the state $\hat{s}_{t}$ and action $\hat{a}_{t}$ taken. The initial continuous state is observed $\hat{z}_{t}$ and it is expected that the final state cost $W$ should deliver a state $z_{t+1}$ that maps on to $\bar{s}_{t+1}$.
The problem is solved in a standard rolling-horizon MPC fashion, that is, sequentially wherein after the first time the solution is computed, the optimal control decision for the first time instant is performed and the state is allowed to evolve, and when the resulting intermediate state has been observed, the process is repeated with the new horizon now shorter by one time step.
We now proceed to describe the components of this operation formally. Real time optimization is performed in the fast time scale, $\{\tau(0),\tau(1),\cdots,\tau(H-1)\}:=\{t,t+h,t+2h,\cdots,t+1-h\}$ with the total number of time steps $H=h^{-1}$.
We definite parameters for constructing the objective and constraint functions in the MPC problem definition as depending on the action to be
$\left\{p^{w}(a_{t}),p^{g}(a_{t}),p^{f}(a_{t}),p^{h}(a_{t}),p^{a}(a_{t})\right\}$
The initial physical state $\hat{z}_{t}$ is observed and the defined optimal control problem is solved below. Note that we can also prescribe a final state set if we want a harder enforcement of landing in the continuous state region corresponding to discrete state space element $\bar{s}_{t+1}$.
Additional details, such as providing warm start solution estimates based on historical planning actions and previous OCP invocations, as well as learning components of the dynamics, are left to a future work on implementation.
At time $\tau_{0}=t=\hat{t}$, the first optimal control problem is solved:
$\min_{z\in\mathbb{R}^{Hd_{z}},v\in\mathbb{R}^{Hd_{v}}}$ $\quad W(z_{(t+1)};p^{w}(a_{t}))+\sum_{\tau(k)=(\tau(0)=t)}^{\tau(H)=t+1-h}\,G\left(z_{\tau(k)},v_{\tau(k)};p^{g}(a_{t})\right)\,,$ (53a)
$\mathrm{s.t.}$ $\quad z_{\tau(k+1)}=f\left(z_{\tau(k)},v_{\tau(k)};p^{f}(a_{t})\right),$ (53b)
$\quad h\left(z_{\tau(k)},v_{\tau(k)};p^{h}(a_{t})\right)\leq 0,\quad v_{\tau(k)}\in\tilde{A}(p^{a}(a_{t}))\,,\quad$ (53c)
$\quad\quad\forall k\in\{0,\cdots,H-1\}$ (53d)
$z_{\tau(0)=t}=\hat{z}_{t}$ (53e)
After the optimization problem is solved to compute $\left\{z^{*}_{t,\tau(:)}(\hat{a}_{t},\hat{z}_{t}),v^{*}_{t,\tau(:)}(\hat{a}_{t},\hat{z}_{t})\right\}$, the first control $\hat{v}(t)=v^{*}_{t,\tau(0)}$ is implemented by the robot controller.
Subsequently, the system is evolved until time $t+h$, and the resulting state $\hat{z}(t-1+h)$ is measured. This is expected to be close to the computed estimated state from the optimization solution at $t$, that is $z^{*}_{t,\tau(1)}$, however, due to real-world state-system mismatch, they are not exactly equal. Still, their proximity enables fast Newton-based methods for computing warm-started solutions. Thus, another optimization problem is solved in real-time. The second problem, solved at and for time $t+h$, is now defined as:
$\min_{z\in\mathbb{R}^{(H-1)d_{z}},v\in\mathbb{R}^{(H-1)d_{v}}}$ $\quad W(z_{(t+1)};p^{w}(a_{t}))+\sum_{\tau=t+h}^{\tau(k)=(\tau(1)=t+1-h)}\,G\left(z_{\tau(k)},v_{\tau(k)};p^{g}(a_{t})\right)$
$\mathrm{s.t.}$ $\quad z_{\tau(k+1)}=f\left(z_{\tau(k)},v_{\tau(k)};p^{f}(a_{t})\right),$
$\quad h\left(z_{\tau(k)},v_{\tau(k)};p^{h}(a_{t})\right)\leq 0,\quad v_{\tau(k)}\in\tilde{A}(p^{a}(a_{t})),$
$\quad\quad\forall k\in\{1,\cdots,H-1\}$
$z_{t}=\hat{z}(\tau(1))$
and again $\hat{v}(t+h)=v^{*}_{t+h,\tau(1)}$ is implemented and state $\hat{z}(\tau(2))=\hat{z}(t+2h)$ is observed. This is repeated for $\tau(k)\in\{t-1+h,\cdots,t-2h,t-h\}$ in this rolling horizon to finally obtain the complete realization,
$\left\{\left\{\hat{z}(\tau),\hat{v}(\tau)\right\},\,\tau=t,\cdots,t+1-h,\right\}\cup\hat{z}(t+1)$
In addition, a real-time reward $\hat{R}^{e}(t+1)$ is observed externally at the conclusion of the problem, together with the evaluation of the action closed loop cost functional $\hat{R}^{o}(t+1)$. This quantity $\hat{R}^{e}(t+1)$ represents mission-relevant information that occurred in the course of the operation. For instance, if, in performing one of the control actions, the robot observed negative emotional affect in the human, this can enter into $\hat{R}^{e}(t+1)$ in a manner so as to inform the upper layer, and ultimately the task can be planned to be performed more agreeably to the human, that is, without action $\hat{a}_{t}$.
6.4 Integrating State representations
To connect the high-level planning state representation, which consists of logical atoms (Sec. 4), with the low-level numeric state vector representation used in control (Sec. 3), we introduce a formalization based on the fuzzy membership functions (Sec. 5). This allows the high-level symbolic representation to be grounded in the continuous, low-level control space while preserving interpretability.
With the high-level state defined in terms of fuzzy logic values, we can attempt an integrated planning and control execution cycle. This consists of (i) selecting a high-level action $a_{t}$ by the planner based on current state $s_{t}$, (ii) executing the corresponding low-level optimal control process yielding $z_{t+1}$, and (iii) mapping back to a high-level symbolic description $s_{t+1}$ from the observed low-level system state $z_{t+1}$. The process then continues analogously with the next action $a_{t+1}$.
The high-level state $s$ is composed of a set of logical propositions (atoms) $\{L_{1},L_{2},\dots,L_{M}\}$ that describe discrete properties of the system, such as “robot A is at place B” (At(A,B)) or “object C is grasped by A” (grasped(C,A)). These logical atoms must be somehow inferred from the underlying low-level state $z$, which consists of continuous physical quantities like position, velocity, or force. To achieve this, each possible logical atom $L_{m}$ is associated222possibly associated on the level of predicates from the domain language $\mathcal{L}$ defining $s_{t}$ (Sec. 4.2.1), i.e. all atoms $L_{m}$ grounded from the same predicate $P(L_{m})$ share the same membership function. with a fuzzy membership function $\mu_{L_{m}}(z)$ that maps a low-level state to a truth value in $[0,1]$:
$\mu_{L_{m}}(z):\mathbb{R}^{d_{z}}\to[0,1].$ (54)
Here, $\mu_{L_{m}}(z)=1$ indicates that the low-level state fully satisfies $L_{k}$, while $\mu_{L_{m}}(z)=0$ indicates complete disagreement, with intermediate values capturing partial satisfaction.
For instance, if $L_{1}$ corresponds to the atom At(RobotA,LocationB), its membership function could be defined as
$\begin{array}[]{l}\mu_{At}(z)=\mathbf{1}\left(\|p_{\text{RobotA}}-p_{\text{LocationB}}\|^{2}\leq\bar{d}^{0}_{At}\right)\\ \qquad+\mathbf{1}\left(\bar{d}^{0}_{At}\leq\|p_{\text{RobotA}}-p_{\text{LocationB}}\|^{2}\leq\bar{d}^{1}_{At}\right)\exp\left(-\frac{\|p_{\text{RobotA}}-p_{\text{LocationB}}\|^{2}-\bar{d}^{0}_{At}}{\sigma^{2}}\right),\end{array}$ (55)
where $p_{\text{RobotA}}$ and $p_{\text{LocationB}}$ are their respective positions, and $\sigma$ controls the smoothness of the transition. Note the truncation of the distribution: if the robot is within distance $\bar{d}^{0}_{At}$ of LocationB, we declare that it is At that location with certainty, and if it is a distance at least $\bar{d}^{1}_{At}$ away, then he is crisply declared to not be at the location, with a fuzzy transition function in the annulus defined by the second indicator in the expression.
The high-level state $s$ can then be represented as an operation on a vector of fuzzy truth values:
$s^{\mu}=T^{s}(\mu_{L_{1}}(z),\mu_{L_{2}}(z),\dots,\mu_{L_{m}}(z);a_{t},s_{t}).$ (56)
ensuring a continuous mapping between low-level numeric and high-level symbolic descriptions, allowing logical planning (reasoning) to be performed in a way that adapts to the underlying control dynamics.
6.5 Updating the Planning State and Model
After execution of the low-level control, the system reaches some observed state $\hat{z}(t+1)$. The high-level fuzzy state then corresponds to some new fuzzy truth values:
$\hat{s}^{\mu}_{t+1}=T^{s}(\mu_{L_{1}}(\hat{z}(t+1)),\dots,\mu_{L_{M}}(\hat{z}(t+1));a_{t},s_{t}).$ (57)
where $\hat{s}^{\mu}$ is now a fuzzy realization of the state observed and $T^{s}:[0,1]^{M}\times\mathcal{A}\times\mathcal{S}\to[0,1]$
The actual state $\hat{s}_{t+1}$ can now be defined as an appropriate defuzzification operation $\mathcal{D}^{\iota_{t+1}}:[0,1]\to\{0,1\}$ where $\iota_{t+1}\in\mathcal{I}_{\mathcal{D}}$ is one of a set of possible operations of size $\left|\mathcal{I}_{\mathcal{D}}\right|=I_{D}$.
A principled, albeit computationally extensive, approach to identify the actual high-level state would be to select the state that best matches the observation:
$\hat{s}_{t+1}=$ $\mathcal{D}^{\iota_{t+1}}\left(T^{s}(\mu_{L_{1}}(\hat{z}(t+1)),\dots,\mu_{L_{M}}(\hat{z}(t+1));a_{t},s_{t})\right)$
$=\arg\max_{s\in\mathcal{S}}\bigwedge_{k}\mu_{L_{m}}(\hat{z}(t+1)).$
Recall that at the initiation of the process, there was an expected state $\bar{s}_{t+1}$. This was defined to be computed as the model estimate:
$\bar{s}_{t+1}=\arg\max_{s\in\mathcal{S}}p_{\Psi}^{S}(t)=\arg\max_{s\in\mathcal{S}}p_{\Psi}^{S}(\theta^{\psi};a_{t},s_{t})$
We can now use the observed state $\hat{s}_{t+1}$ to update the model $\theta^{\psi}$, e.g., by a gradient update with stepsize $\gamma_{t}$,
$\begin{array}[]{l}\theta^{\psi}_{t+1}=(1-\gamma^{\psi}_{t})\theta^{\psi}_{t}+\gamma^{\psi}_{t}\frac{\partial\mathbb{P}_{p_{\Psi}^{S}(t+1)}}{\partial\theta^{\psi}}(\hat{s}_{t+1})-\frac{\gamma^{\psi}_{t}}{S-1}\frac{\partial\mathbb{P}_{p_{\Psi}^{S}(t+1)}}{\partial\theta^{\psi}}(\bar{s}_{t+1})\\ \qquad\qquad-\frac{\gamma^{\psi}_{t}}{(S-1)^{2}}\sum\limits_{s\in\mathcal{S}\setminus\hat{s}_{t+1}}\frac{\partial\mathbb{P}_{p_{\Psi}^{S}(t+1)}}{\partial\theta^{\psi}}(s)\end{array}$ (58)
Updating the Upper Level Model
Consider that a reward $\hat{R}_{t+1}$ is observed upon the realization $(\hat{a}_{t},\{\hat{z}_{\tau},\hat{v}_{\tau}\},\hat{s}_{t+1})$. With a robotic aide, this could be defined by either prompting the human for direct feedback, observing the human’s expression and estimating emotional affect, as well as externalities on the surrounding environment judged and remarked upon by observers, e.g., in a human-feedback pre-training operation of the system.
$\begin{array}[]{l}\hat{R}(t+1)=R(\hat{R}^{e}(t+1),\hat{R}^{o}(t+1))\\ \hat{R}^{o}(t+1)=W(\hat{z}(t+1);p^{w}(a_{t}))+\sum\limits_{\tau(k)=(\tau(0)=t)}^{\tau(H)=t+1-h}G(\hat{z}(\tau(k)),\hat{v}(\tau(k));p^{g}(a_{t}))\end{array}$
If the computed state $\hat{s}_{t+1}\notin\mathcal{G}$, then we use the intermediate reward to learn the intermediate cost function estimate, say by:
$\theta^{C}_{t+1}=(1-\gamma^{C}_{t})\Phi^{C}_{t+1}(\theta^{C}_{t},s_{t},a_{t})+\frac{\partial\Phi^{C}(\theta^{C}_{t};a_{t},s_{t})}{\partial\theta^{C}}(\hat{R}_{t+1}-\bar{C}_{t+1})$ (59)
and if the state is in the goal set $\hat{s}_{t+1}\in\mathcal{G}$, then a similar update is made to $\bar{C}_{T}$. However, in addition, the loop ceases and the robot is idle until the next task arrives.
The action-value function can be updated in standard fashion:
$\theta^{V}_{t+1}=(1-\gamma^{V}_{t})\theta^{V}_{t+1}V(s_{t},a_{t})+\frac{\partial\Phi^{V}(\theta^{V}_{t};a_{t},s_{t})}{\partial\theta^{V}}(\hat{R}_{t+1}-\bar{V}_{t+1})$ (60)
where now $\bar{V}_{t+1}$ is the action value function estimated by $\bar{V}_{t+1}=\Phi^{V}(\theta_{t}^{V};\hat{a}_{t},\hat{s}_{t})$. Observe that the difference ensures the right sign is used to update the parameters relative to the gradient.
6.6 Fuzzy control relaxation alternative
Alternatively, instead of explicitly identifying a discrete high-level state $s_{t+1}$ after observing $\hat{z}(t+1)$, one may employ a fuzzy control approach (Sec. 5), where the action $a_{t+1}$ selection is performed directly in the low-level continuous space. In this case, the high-level planner can be replaced by a fuzzy controller $\tilde{\zeta}$ that directly determines the next action based on the computed fuzzy memberships:
$\hat{a}_{t+1}=\hat{a}^{\zeta}_{t+1}\tilde{\zeta}(\mu_{L_{1}}(\hat{z}(t+1)),\dots,\mu_{L_{M}}(\hat{z}(t+1)).$ (61)
This formulation removes the need for explicit high-level symbolic state inference by treating the fuzzy memberships as the effective state representation for decision-making. The mapping $\tilde{\zeta}$ may be implemented akin to a Mamdani fuzzy rule-based system (Sec. 5), for example:
IF $\begin{array}[]{l}\text{IF }\mu_{L_{1}}\text{ is high AND }\mu_{L_{2}}\text{ is low, THEN }a_{t+1}=a_{1},\\ \text{IF }\mu_{L_{2}}\text{ is high, THEN }a_{t+1}=a_{2}.\end{array}$ (62)
To relax even further, instead of selecting an intermediate high-level action $a_{t}$, the fuzzy inference system could even directly output the low-level control $v(\tau)$. This would further eliminate any need for the high-level discrete description, as well as the MPC procedure
$v(\tau)=\tilde{\zeta}(\mu_{L_{1}}(\hat{z}(\tau)),\dots,\mu_{L_{m}}(\hat{z}(\tau))).$ (63)
at the expense of interpretability and optimality.
Hybrid Approach
Ultimately, the policy can be defined as some hybrid approach with respect to the computed policy:
$(\hat{a}_{t},v_{t})\sim\Delta\left(\pi;\hat{a}^{g}_{t},\hat{a}^{\epsilon}_{t},\hat{a}^{\sigma}_{t},\hat{a}^{c}_{t},\tilde{\zeta}_{t}\right)$
where $\Delta$ is the unit simplex over the arguments. That is, the Algorithm can compute a classical planning algorithm solution, an RL policy, or a fuzzy control relaxation scheme with some probability across the options.
6.7 Scheduling-Planning Integration
We propose that the apex problem is one of Scheduling. Namely, within each daytime period of 16 hours, the robot has a series of tasks that it can perform for the human it is of service to: $\mathcal{T}=\{T_{b},T_{l},T_{d},T_{m_{1}},T_{m_{2}},T_{w},T_{e}\}$ corresponding to feeding breakfast, lunch and dinner (i.e., $f(T_{b})<s(T_{l})$), providing medicine regime 1 and 2, and an event triggered task of assistance to the use the toilet and provision of some entertainment (music, jokes, etc.). We consider that the corresponding events follow Poisson distributions with parameters $\lambda_{b}$ and $\lambda_{e}$.
We can consider that the goal is to ensure all tasks are completed within the day, with some regularization towards spreading the tasks evenly throughout the day. We can see that this intuitively yields a solution of, for instance, 30 minutes duration for each task corresponding to arranging to perform $\{T_{b},T_{m_{1}},T_{l},T_{m_{2}},T_{d}\}$ sequentially with about, e.g., $f(T_{b})=s(T_{m_{1}})-2.5$ hours. However, in addition, we add another objective term $\mathbf{F}$ defining the overall quality of the experience of humans throughout their interaction with the agent as a function of the tasks scheduled. We consider that this function is initially unknown, but is a function of the tasks that are scheduled and when, as well as an objective $F_{T}$ corresponding to the positive experience of that given task as it happened in real time. Day by day, $\mathbf{F}$ and its components $\{F_{T}\}$ for $T\in\mathcal{T}$, are learned from interaction with the human, and the Scheduler is optimized for peak human service experience.
If the toilet event is triggered, $T_{w}$ is performed and any other task ceases and is completed following $f(T_{w})$. If the entertainment event is triggered, the task is performed simultaneously with the meals, while it is delayed until after $f(T_{m_{1}})$ and $f(T_{m_{2}})$. These rules can all be encoded into logical constraints. Note that if an event is triggered in a manner that interrupts a scheduled task, then the schedule must be recomputed.
We can consider that there is an event that is triggered by, e.g., erratic behavior, irregular toilet behavior, change in complexion, etc. that notifies an external human health worker.
We can consider that each task spawns a planning problem. That is, for duration $d(T)$, a planning problem $(\mathcal{P}_{T},\mathcal{S}_{T},\mathcal{A}_{T},\Phi_{T})$ (with $\Phi_{T}=(\psi(T),\phi_{G}(T))$ being the dynamics) is executed. The goal to be satisfied in the planning problem is defined by completing the task that it corresponds to, which can be defined to be $n\leq K(t)$, where $K(t)$ is the maximum number of states that can be traversed in the time duration $d(t)$ defined for task $t$ and $s_{n}\models\phi_{G}$.
As an example, for $T_{l}$, actions could include moving to the kitchen, which must happen first, slicing vegetables, boiling stew, sprinkling a teaspoon of dill, etc. Some of these, in sequence, construct a meal. The previous Section defined the operation of the planning problem that sequences the set of actions to perform the triggered task. These are also associated with a lower level control system. Note that while the physical dynamics of the robot, that is, $f$ above, exhibit significant invariance across Tasks, and thus planning problems, the constraints $h$ as well as the terminal and intermediate costs $W$ and $G$ are expected to be defined in a manner influenced by the valence of the goal(s) associated to the Task.
There is an additional observed reward $\hat{R}(T_{l})$ that quantifies how pleasant the experience was for the human. This can be measured by, e.g. a 1-10 rating given by the human after each task and/or reading emotional affect during the operation of the planning task. In addition a live or delayed video recording observed by a health expert can grade the performance in optimality to patient physical and mental health  [^3] We expect that this human-in-the-loop pretraining will be important for adequate alignment to the operation’s needs and goals.. This reward is meant to learn an additional Value Function for the choice of Task, but must be learned through interaction with the environment. Recall that for each task/planning problem, there are multiple sequences of actions that yield the goal state, and the goal is to find the one that maximizes the objective $F^{T_{l}}(s,a;T_{l},\mathcal{E}(s_{0}))$. The incorporation of an additional parametrization by vector $\mathcal{E}$ represents the time of day the task is performed as well as other possibly relevant covariates and and a vector describing the read emotional affect of the human. Then the Task Scheduler can learn the Value of a particular planning problem. Note that it depends on the initial state, and is otherwise determined by the algorithm defining vectors $\{s_{t},a_{t}\}_{t\in[n]}$.
7 Conclusion
The increasing interest of deploying AI in real world scenarios, from autonomous vehicles and drones to robotics in manufacturing and care, presents the necessity of enforcing reliability through first principles domain knowledge. Increasingly, there is a complex layered taxonomy of tasks and behaviors, and interpretability is essential for fostering trust and robust performance. This paper presents a first attempt at simultaneously integrating classical planning, which encodes crisp goals in a solid logical framework, through to MPC, the most state of the art the technology for fast real time control, while incorporating classic RL tools for learning the rewards from the environment. We defined the entire end-to-end Scheduling to Planning to Control, including both the operation and the co-learning across layers. We hope the work opens up a productive research program towards the management and control of autonomous agents.